### v2

# Point2Primitive: CAD Reconstruction from Point Cloud by Direct Primitive Prediction

Abstract
Recovering CAD models from point clouds, especially the sketch-extrusion process, can be seen as the process of rebuilding the topology and extrusion primitives. Previous methods utilize implicit fields for sketch representation, leading to shape reconstruction of curved edges. In this paper, we proposed a CAD reconstruction network that produces editable CAD models from input point clouds (Point2Primitive) by directly predicting every element of the extrusion primitives. Point2Primitive can directly detect and predict sketch curves (type and parameter) from point clouds based on an improved transformer. The sketch curve parameters are formulated as position queries and optimized in an autoregressive way, leading to high parameter accuracy. The topology is rebuilt by extrusion segmentation, and each extrusion parameter (sketch and extrusion operation) is recovered by combining the predicted curves and the computed extrusion operation. Extensive experiments demonstrate that our method is superior in primitive prediction accuracy and CAD reconstruction. The reconstructed shapes are of high geometrical fidelity.
Figure 1: CAD reconstruction pipeline including (a) The Point2Primitive network first segmented the input point cloud into clusters belonging to the same extrusion and corresponding barrel points, and the curves of each extrusion primitive are directly learned and predicted. CAD reconstruction can be achieved by combining the reconstructed extrusion sketch and extrusion operations computed from the barrel points, which is easy to edit in downstream tasks. (b) The illustrations of the sketch-extrude process, extrusion segmentation, and extrusion primitives. The topology describes how extrusion primitives are interrelated.
**footnotetext: Corresponding Author
1 Introduction
CAD reconstruction from point clouds is a longstanding objective [10]. In CAD, reverse engineering encompasses techniques that extract CAD models from existing products by transforming unstructured data (e.g. point clouds, meshes, or printed schematics) into structured representations of 2D or 3D geometries. The sketch-and-extrude process is one of the most fundamental CAD modeling commands, offering a highly intuitive and versatile way to generate a broad array of shapes, as demonstrated in [36], and investigating the reverse of this process is valuable to the community.
However, the unordered and unstructured nature of point clouds poses significant challenges for high-level manipulation and efficient editing of their underlying geometries [27]. Some methods try to solve this problem and reconstruct the shapes through optimization methods [19, 1, 30, 15, 18, 26]. The recent availability of significant boundary representation (B-Rep) datasets [36, 9, 31] has introduced data-driven frameworks to reverse engineering in the CAD world. Numerous neural networks have been developed to segment point clouds into clusters to represent distinct surface regions [13, 32, 11, 21]. These segmented points are then fitted to surface primitives (either parametric or freeform), allowing the topological assembly of these primitives to reconstruct the CAD models. In contrast, we focus on recovering the CAD models that describe the sketch-and-extrude modeling process. To this end, we employ a point segmentation network to recover topological structures (extrusion primitives) by segmenting the point clouds into clusters associated with individual extrusion primitives. By predicting the curves and extrusion operation of each extrusion primitive, the CAD models can be reconstructed.
Specifically, the sketch serves as the foundation of parametric CAD modeling, with the quality of the final CAD model depending critically on the precision of its sketches [5]. In the parametric CAD workflow, close curves constitute the sketch, and 3D models are then constructed by extruding the sketches and grouping these extrusions into structured assemblies. Recently, implicit representations of 3D shapes [35, 23], have inspired new approaches that employ SDFs to represent sketches, enabling sketch and CAD reconstruction through SDF encodings [34, 29, 14]. However, decoding SDF embeddings into sketches typically results in CAD models with unintended curved edges and the reconstructed sketch is not easy to be directly edited. Inspired by large language models (LLMs) [4], recent works [37, 39, 22] propose to generate CAD modeling sequences using language-based models. However, our findings indicate that LLMs struggle to implicitly learn hierarchical representations of topology (e.g. CSG structures) [41] using only reconstruction loss, such as chamfer distance and primitive loss, which can result in limited generalization. SkexGen [38] strives to address this challenge by introducing separate sequence encoders for modeling topology and geometry independently to facilitate related design generation. In response, we propose representing sketches through their fundamental curves and explicitly recovering their extrusion primitives [41] by decomposing input points into distinct extrusion sets.
This paper presents a novel approach for reconstructing sketch-and-extrude CAD models using an end-to-end point-to-primitive (Point2Primitive) network. Specifically, our method first segments the input point cloud into distinct extrusion primitives and corresponding barrel/base points. Each extrusion sketch is then predicted by directly predicting the type and parameter of its curves through an improved transformer network. By combining the predicted curves and the extrusion operation calculated from the barrel points, the CAD models can be rebuilt. Using curves as sketch representation, our approach reconstructs fundamental shapes in sketches (such as lines, arcs, and circles), resulting in accurate, geometry-preserving, and easily editable 3D CAD models. Our primary contributions are summarized as follows:
  • We propose a CAD reconstruction method based on the Point2Primitive network, which leverages an enhanced transformer to predict sketch and PointNet++ for extrusion segmentation. By representing the sketch through its fundamental curves rather than using SDF, our approach achieves higher geometric fidelity.
  • We introduce an improved transformer decoder along with a sketch hierarchy representation. The curve parameters are encoded as position embeddings to guide the attention weight in locating the target. The parameters are directly predicted and optimized in an auto-regressive way, leading to high accuracy.
  • We present a complete data generation pipeline that can produce extrusion segmentation labels, barrel/base segmentation labels, and sketch primitives from the CAD modeling sequence saved in JSON files, which can be used for future research.
  • We conduct a comprehensive experiment with several CAD reconstruction results to analyze the proposed method thoroughly. Also, we reproduced the data generation methods used by other methods for comparison. An ablation study is conducted to fully investigate the proposed method.
2 Related works
2.1 CAD Reconstruction via Primitive Detection
Several learning-based approaches have also been proposed to fit geometric primitives to point clouds. These methods first segment the input points into clusters corresponding to the same surface primitives and fit the clustered points to the parametric surface primitives [13, 32, 11, 17] or free surfaces [21]. Except for surface primitives, UCSG-Net[7] and CSG-stump [28] learn the CSG parse tree that interrelates the predicted geometry primitives to reconstruct shapes. The reconstructed CAD models are in B-rep format, which is not as easy to edit as the sketch-and-extrude modeling procedure. In this paper, we explore the inverse of the sketch-and-extrude process and strive to recover the CAD modeling procedure from the point cloud. Therefore, we segment the point clouds into clusters belonging to the same extrusion primitive and explicitly reconstruct every element of the extrusion primitives. Instead of fitting points to the parametric or free surface, we employ an improved transformer to learn and predict the curves directly.
2.2 Sketch-and-extrude CAD Reconstruction
Some new solutions use SDF to represent sketches following the recent implicit 3D shape representations [23, 35]. ExtrudeNet [27] and SECAD-Net [14] can learn implicit sketches and differentiable extrusions. Point2Cyl reconstructs 3D shapes through sketch regression supervised by the latent embeddings of its SDF. However, these methods utilize the implicit fields for sketch representation, leading to curved edges of the reconstructed shapes. Meanwhile, the reconstructed sketch is hard to be directly edited because it is represented by the SDF. Further transformation from SDF to sketch further brings in extra parameter errors. In this paper, we propose to represent a sketch using its curves and directly learn and predict the curve types and parameters from the point clouds in an object detection manner [33]. To this end, we propose the center-prior primitive definition (detailed in Section 3) and the improved transformer decoder (detailed in Section 4.4) to achieve precise and editable sketch reconstruction, which makes the final CAD reconstruction easy to be edited. Also, the most related work to us is the Point2Cyl, and we conduct an extra comparison as briefed in Section .
2.3 CAD Generation
Following Large Language Model (LLM) [4], [37] is the first to explore the CAD sequence generation of 3D shapes based on VAE [8]. [39] utilizes a vector quantized VAE to generate modeling sequences. [22] propose to generate CAD models step-by-step based on multimodal diffusion. It should be noted that the CSG structure [41] is broken when tokenizing the CAD modeling sequence. We find that the implicit representation learning of the topology is challenging for LLM if only supervised by geometry loss (such as Chamfer Distance) or command accuracy. SkexGen [38] propose to use two sequence encoders separately for topology and geometry learning. However, CAD generation methods leveraging VAE [8] structure and language model decoder still limit its performance on reconstruction tasks because of poor generalization. Also, language models struggle with numerical representation and mathematics [40], causing inaccuracies in command parameters and limited performance beyond the training dataset (detailed in Section 5.4). In this paper, we present a CAD reconstruction method based on the Point2Primitive network. We explicitly recover the topology through point segmentation and directly predict the sketch curves and extrusion operation. By explicit precise primitive parameter prediction, the CAD reconstructions are of high geometry fidelity and good generalization.
3 Preliminaries
In CAD terminology, the sketch consists of multiple closed curves $S_{i}$. There are three types of curves: line $\mathit{L}$, arc $\mathit{A}$, and circle $\mathit{R}$ in our sketch hierarchy. Therefore, the sketch $S_{i}=[c_{1}^{i},...,c_{N}^{i}]$, where $c_{j}^{i}\in(\mathit{L},\mathit{A},\mathit{R})$ for all $j\in[1,...,N]$, can be seen as a set of curves. The goal of predicting sketch is to train the parameters $\theta$ of the Point2Primitive network $f_{\theta}$ on the training pair $({p_{i}},\hat{S}_{i})$ to make a good approximator of the set prediction as follows:
$f_{\theta}\left(\{p_{k}\mid p_{k}\in P_{\mathbb{E}_{i}}\}\right)\approx\left\{c_{1}^{i},...,c_{j}^{i}\mid c_{j}^{i}\in(\mathit{L},\mathit{A},\mathit{C})\right\}$ (1)
where $P_{\mathbb{E}_{i}}\in R^{N_{E_{i}}\times{3}}$ is the points of the $i\mbox{-}th$ extrusion $\mathbb{E}_{i}$.
Curve Type Parameters
$L$ $(Line)$ $x_{m}$ $y_{m}$ $x_{1}$ $y_{1}$1 -1
$A$ $(Arc)$ $x_{m}$ $y_{m}$ $x_{1}$ $y_{1}$ $x_{2}$ $y_{2}$
$R$ $(Circle)$ $x_{c}$ $y_{c}$ $r$1 -1 -1
Table 1: Curve definition. We provide the types and vector representations of curves used in this work. Unused parameters in curve vector are padded with $-1$.
To work with the improved transformer decoder, we formulate the definition of curve parameter in a centra-prior way, as shown in Table 1. Each primitive parameter is primarily presented by its center.
Line is represented with its midpoint ($x_{m},y_{m}$) and another endpoint ($x_{1},y_{1}$). Arc is defined with its midpoint ($x_{m},y_{m}$), start points ($x_{1},y_{1}$) and endpoints ($x_{2},y_{2}$). The circle parameter consists of the center point ($x_{c},y_{c}$) and radius $r$. Each curve $C_{i}=(t_{i},\rho_{i})$ is defined by its primitive types $t_{i}\in{\mathbb{R}}$ and primitive parameters $\rho_{i}\in\mathbb{R}^{1\times{6}}$ with unused parameters are padded with $-1$. Therefore, the first two elements of the parameters $\rho_{i}$ can be encoded as position embeddings to guide the attention in the improved transformer decoder as briefed in Section 4.4
In the sketch-extrude process, a sketch $S_{i}$ is extruded by an operation $E_{i}$ to form an extrusion $\mathbb{E}_{i}=\left[S_{i},E_{i}\right]$, and multiple such extrusions collectively form a solid shape $\mathbb{S}=\{\mathbb{E}_{1},...,\mathbb{E}_{K}\}$. The shape $\mathbb{S}$, therefore, can be represented by the SDF defined by the collection of the extrusions. In our method, the $\{\mathbb{E}_{1},...,\mathbb{E}_{i}\}$ is reconstructed via point segmentation, while the sketch $S_{i}$ is predicted through an improved transformer network.
4 Point2Primtive
4.1 Overview
We propose Point2Primitive, a deep model for CAD reconstruction from point clouds. As illustrated in Figure 2, our method utilizes a point segmentation network to cluster points that belong to the same extrusion. Base/barrel points are also classified during extrusion segmentation. Then an improved transformer is introduced to predict the curve type and parameters of each extrusion. As defined in Equation 1, curve prediction can be seen as a set prediction problem. Therefore, we use the transformer network to solve the set-to-set problem and make some improvements to the transformer decoder, as shown in Figure 2. The transformer encoder is used to refine the barrel point features from the segmentation (Seg) head. The Seg head is detailed in Appendix 4.2, where the barrel point encoder is developed based on the PointCNN [16] with fixed grid points as representation points instead of random points.
4.2 CAD Dataset Generation
In this paper, we propose to generate the training dataset based on the Signed Distance Field (SDF). An SDF $SDF(p_{eval},b)$ is a continuous function that, for a given spatial eval point $p_{eval_{i}}$, outputs the eval point’s distance to the closest boundary defined by $b$, whose sign denotes whether the point is inside (negative) or outside (positive) of the watertight surface. The data generation pipeline can convert any CAD sequence saved in JSON files following the style of the Fusion360 Gallery dataset [36] to the training data pairs.
Given an extrusion point set $P_{\mathbb{E}_{i}}\in R^{N_{\mathbb{E}_{i}}\times{3}}$ and its sketch $S_{i}$. We first project the points $P_{\mathbb{E}_{i}}$ to the sketch plane as eval points according to the transform matrix $W_{i}$ defined by its extrusion axis $e_{i}$. Then, its SDF is calculated to the boundary defined by the sketch $S_{i}$. The batch SDF calculation will be detailed in the Appendix 5.1. The barrel label $b_{i}$ of each point can be defined as follows:
$b_{i}=\text{bool}(\text{SDF}(W_{i}\cdot{P_{\mathbb{E}_{i}}},S_{i})\leq\text{thresh})$ (2)
where $thresh$ controls the label precision and $\text{thresh}=0.01$ in this paper.
As for the Extrusion Label, each extrusion mesh is first built using pythonOCC [24] by the CAD sequence. Then, the SDF of the input points to each extrusion boundary defined by its mesh is calculated. The extrusion label of each point is set as its closest extrusion mesh.
4.3 Extrusion Segmentation
Our method utilizes a point segmentation network to cluster points and rebuild the topology, as shown in Figure 2. The Point2Primitive is insensitive to the implementation of the point segmentation network. We choose the PointNet++ [25] as the point encoder and the decoder, which is detailed in Appendix 4.1.
Figure 2: Overview of Point2Primitive. Our model consists of two main parts: extrusion segmentation and sketch prediction. In extrusion segmentation, a PointNet++ model is used to segment the barrel points and extrusions from the given point cloud to rebuild the topology. In sketch prediction, the point features from the Seg head are first refined using a vanilla transformer encoder and decoded into curves through the improved transformer decoder.
4.4 Improved Transformer Decoder
To find the implicit design features encoded in the point features, we propose an improved transformer decoder to convert the refined point features into curves and directly predict the parameters. As shown in Figure 3, the curve embeddings encode the type information, and the curve parameters are formulated as positional queries. By decomposing the learnable queries into type and parameter embeddings, explicit geometric priors are introduced, avoiding the abstract learning process and leading to high accuracy.
Figure 3: The improved Transformer decoder. The curve parameters are formulated as the position queries, while the curve embeddings encode the curve type. The curve parameters are updated in each Transformer layer through the predicted $\delta$.
Parameter as Position Encoding.
The curve parameters are formulated in a center-prior format detailed in Section 3, so the curve parameters can be formulated as position encoding $PE_{i}$ to guide the attention. Given the curve parameters $\rho^{i}_{j}$ as the parameters of the $j\mbox{-}th$ curve in the $i\mbox{-}th$ sketch $S_{i}$. The corresponding positional queries $PE_{j}^{i}$ are generated by:
$PE_{j}^{i}=\text{MLP}(f_{pe}(\rho^{i}_{j}))=\text{MLP}(f_{pe}(x_{m_{j}}^{i},y_{m_{j}}^{i}))$ (3)
where $f_{pe}$ is the function that generates the sinusoidal embeddings from the curve center. Multi-layer perception (MLP) and ReLU activations are used to produce the position encodings $PE_{j}^{i}$.
Direct Curve Parameter Prediction.
The curve prediction procedure can be seen as feeding the positional queries (curve parameters) and curve embeddings (type queries) into the decoder to probe the curves $c^{i}_{j}$ that correspond to the position encodings while having similar patterns with the content in the point features. Furthermore, we propose to predict the curve parameters directly. The curve parameters are updated layer-by-layer in an autoregressive way to fit the target, as shown in Figure 3. The layer-by-layer update is based on the decoder output as follows:
$\delta x_{m_{j}}^{i},\delta y_{m_{j}}^{i},\delta p_{2_{j}}^{i},\delta p_{3_{j}}^{i},\delta p_{4_{j}}^{i},\delta p_{5_{j}}^{i}=\text{MLP}(O_{n})$ (4)
where $O_{n}$ is the output of the $n\mbox{-}th$ decoder layer. The curve parameters are updated to $[x_{m_{j}}^{i^{\prime}},y_{m_{j}}^{i^{\prime}},p_{2_{j}}^{i^{\prime}},p_{3_{j}}^{i^{\prime}},p_{4_{j}}^{i^{\prime}},p_{5_{j}}^{i^{\prime}}]$ in each layer, as shown in Figure 3. All intermediate parameter outputs are supervised by ground truth. DeNoising [12] is utilized to accelerate the convergence.
4.5 Loss Function
The Point2Primitive is trained using a multi-task, non-convex objective composed of extrusion segmentation loss $\mathcal{L}_{\mathbb{E}}$, base-barrel classification loss $\mathcal{L}_{{bb}}$, normal $\mathcal{L}_{norm}$ and sketch $\mathcal{L}_{skh}$ prediction losses:
$\mathcal{L}=\mathcal{L}_{\mathbb{E}}+\mathcal{L}_{{bb}}+\mathcal{L}_{norm}+\mathcal{L}_{skh}$ (5)
where the cross entropy loss is used in $\mathcal{L}_{\mathbb{E}}$ and $\mathcal{L}_{{bb}}$ while L-1 loss is used in $\mathcal{L}_{norm}$. The sketch prediction loss $\mathcal{L}_{skh}$ can be formulated as:
$\mathcal{L}_{skh}$ $=\sum\limits_{i=1}^{N_{C}}l_{focal}(\hat{t}_{i},t_{i})+\beta\sum\limits_{i=1}^{N_{C}}\sum\limits_{j=1}^{N_{p}}\mathcal{L}_{p}(\hat{\rho}_{i,j},\rho_{i,j})$ (6)
$\mathcal{L}_{p}$ $=L_{1}(\hat{\rho}_{i,j},\rho_{i,j})+L_{2}(\hat{\rho}_{i,j},\rho_{i,j})$
where $l_{focal}(\cdot,\cdot)$ represents the focal loss [20] and $L_{1}(\cdot,\cdot)$ and $L_{2}(\cdot,\cdot)$ represents the L1 and L2 loss, respectively. Both Denoising Loss and Curve Loss use the sketch prediction loss $\mathcal{L}_{skh}$, as shown in Figure 2. More specifically, in the L2 loss, we randomly select $n$ points on each curve and compute the chamfer distance as the L2 loss. $N_{P}$ is the number of parameters ($N_{p}=6$ is this paper) while $N_{C}=30$ is the number of curves. $\beta$ is the weight to balance both terms ($\beta=2$ in this paper) and $\hat{\cdot}$ denotes the ground truth value.
5 Experiments
5.1 Experimental Setups
Datasets. To evaluate the proposed model, we conduct experiments on DeepCAD [37] and Fusion 360 Gallery [36]. The training data is generated as described in Sec. 4.2.
Implementation Details. We use PointNet++[25] for extrusion segmentation. The Point2Primitive is trained for 400 epochs with a total batch size of 64 and learning rate (lr) $51e-$ 5 with linear warmup and step reduce lr scheduler. As for the DeNoising settings, the noise rate for the curve type and parameters is $0.2$ and $0.3$, respectively, with 5 DeNoising groups. The transformer encoder and decoder contain 6 and 8 transformer layers, respectively, with eight attention heads and latent model dimension 256.
Evaluation Metrics. We adopt command type accuracy ($Acc^{SKH}_{t}$) and parameter accuracy ($Acc^{SKH}_{p}$) for quantitative evaluations of the sketch prediction following [5]. For networks that represent sketch by the implicit nueral field, the tools for convert sketch SDF to command are developed based on the tool by SECAD [14]. Furthermore, following [3], we also report chamfer distance ($CD$), edge chamfer distance ($ECD$), normal consistency ($NC$), and the number of generated primitives ($\Delta\#P$) to measure the quality of the recovered 3D geometry, more details of the quantitative metrics are shown in the Appendix 1.
5.2 Main Results
[b] DeepCAD [37] Fusion 360 Gallery [36] Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$ IR #$\#\Delta P$ $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$ IR #$\#\Delta P$ Methods of Representing Sketch with Implicit SDF Encoding ExtrudeNet [29] 34.58% 31.71% 0.379 0.962 0.849 24.19% 34.34 37.81% 34.39% 0.671 0.808 0.809 23.83% 27.47 Point2Cyl [34] 41.37% 39.41% 0.489 1.027 0.819 3.91% 26.18 42.98% 41.16% 0.529 0.769 0.769 3.87% 26.96 SECAD-Net [14] 45.91% 41.37% 0.341 0.868 0.861 8.03% 32.78 46.82% 42.97% 0.449 0.684 0.813 7.82% 28.61 Generation Methods Based on Language Model DeepCAD [37] 82.61% 73.36% 0.898 1.883 0.823 14.12% 5.89 77.31% 69.81% 7.128 8.729 0.719 13.18% 7.47 HNC-CAD [39] 84.31% 76.71% 0.827 1.064 0.846 6.01% 4.43 80.62% 72.19% 4.381 5.571 0.748 5.92% 6.14 Methods Based on Primitive Fitting UCSG-Net [7] - - 1.849 1.174 0.820 - 14.19 - - 0.952 1.277 0.770 - 11.17 CSG-Stump [28] - - 3.031 0.755 0.828 - 19.46 - - 0.781 0.991 0.744 - 13.08 Our Method 96.14% 86.81% 0.312 0.581 0.897 3.71% 4.14 94.17% 83.52% 0.392 0.571 0.839 3.62% 5.17
Table 2: Evaluation results on the test set of the CAD reconstruction dataset. CAD reconstruction and generation methods are selected for comparison.
To demonstrate its effectiveness, we compare the proposed model to multiple existing methods, covering the methods of representing sketch with SDF encoding [14, 29, 34], the generation methods based on language model [37, 39], and methods based on primitive fitting [7, 28]. The direct reconstruction results are visualized for comparison.
Table 2 summarizes the main results, and we can find that sketch curve type and parameter accuracy of the Point2Primitive are much higher than the other methods while the CD metric of the Point2Primitive is lower. The smallest CD and ECD value indicates that the presented CAD reconstruction method achieves better geometry fidelity while preserving accurate and sharp shape edges. This shows that by directly recovering curves of the extrusion primitive, our method can achieve better performance than the other extrusion-segmentation methods that represent the sketch with an implicit field. As for the number of generated primitives, our method is closer to human designs (The smallest #$\#\Delta P$ value). This improvement is due to the explicit fine-grained reconstruction from each extrusion primitive; every parameter of the curves and extrusion operations are all predicted and optimized.
5.3 Visualization Results
Figure 4: Visual comparison between reconstruction results on the DeepCAD (left) and Fusion360 gallery dataset (right).
Figure 4 demonstrates the visualized reconstruction results on the DeepCAD and Fusion 360 Gallery datasets, respectively. It can be seen that the results of our method are more compact and complete, while the edges are sharp and sketches are accurate. This shows that our method can produce CAD reconstruction of high geometrical fidelity. The methods based on the language model can also produce accurate geometry structure, but the error of the parameters in the sequence results in some models of poor compactness. The primitive detection methods (USCG and Stump) fail to predict some holes in the model. Still, the edges are critically more accurate than the ExtrudeNet, whose reconstructions are less complete.
5.4 Comparison on Augmented DeepCAD Dataset
In addition to the experiments on the original DeepCAD and Fusion 360 Gallery datasets, we conducted extra comparisons on the augmented DeepCAD datasets. More specifically, we add random noise to the parameters of the CAD modeling sequence while keeping the types intact. This noise causes the augmented DeepCAD test set to have model structures similar to the original shape but with some modifications to the geometry details. We train each method on the original DeepCAD dataset and test all the methods on the augmented DeepCAD test sets.
Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$$\#\Delta P$
Methods of Representing Sketch with Implicit SDF Encoding
ExtrudeNet [29] 28.39% 25.61% 0.614 1.117 0.776 36.14
Point2Cyl [34] 34.47% 30.25% 0.518 1.065 0.791 27.96
SECAD-Net [14] 36.61% 31.43% 0.437 1.079 0.806 34.18
Generation Method Based on Language Model
DeepCAD [37] 61.41% 43.71% 5.919 6.883 0.708 7.16
HNC-CAD [39] 63.39% 53.29% 6.864 7.064 0.711 6.56
Methods Based on Primitive Fitting
UCSG-Net [7] - - 2.146 1.273 0.797 14.81
CSG-Stump [28] - - 3.681 0.958 0.804 20.16
Our Method 94.67% 83.81% 0.410 0.607 0.819 4.97
Table 3: Evaluation results on the Augmented DeepCAD dataset
Figure 5: Visual comparison between reconstruction results on the augmented DeepCAD dataset.
Table 3 demonstrates the quantitative results of the augmented DeepCAD dataset. We add the original shape (Origin) to the last column in the visualized results as shown in Figure 5. It can be seen that the methods based on the language model see a direct drop in the metrics of the sketch command type and command parameter, leading to a significant increase in the CD metric. Some reconstruction results by the language model preserve similarities to the original shape, leading to false reconstruction. The generated results are similar to the ground truth but with low geometry accuracies. On the other hand, reconstruction methods can still produce results that satisfy geometry fidelity. This implies that the methods based on the language model are less generalizable than the other methods. Also, both the quantitative metrics and the visual results verify that the proposed method can be generalized to reconstruction tasks beyond the train-test dataset.
5.5 Detailed Comparison with Point2Cyl
Figure 6: Comparision with the Point2Cyl: (a) The network architecture of Point2Cyl (left) and our Point2Primitive (right); (b) The reconstructed sketch by Point2Primitive and Point2Cyl; (c) The extrusion segmentation by Point2Primitive and Point2Cyl.
Point2Cyl is the most related method to the proposed Point2Primitive, so we conduct extra comparisons with Point2Cyl. Figure 6 demonstrates the architecture difference between Point2Cyl and the proposed Point2Primitive. Instead of using IGR [6] and PointNet[2] to encode SDF for sketch SDF prediction, an improved transformer is utilized in the Point2Primtive to convert barrel points to curves directly. Thus, the sketch loss is the curve prediction loss rather than the regression loss of the sketch SDF encodings, which introduces more explicit geometry information. Also, we utilize the technique based on the jet fitting [1] (for sketch prediction) and Hough Transform [19] (for extrusion axis) as the optimization methods for comparison.
Metrics Ours Ours Point2Cyl[34] Point2Cyl[34] JF[1]
w/o $\mathcal{L}_{sketch}$ w/o $\mathcal{L}_{sketch}$ HT[19]
Ext. Seg IoU 0.814 0.862 0.801 0.817 0.513
BB. Seg Acc 0.902 0.902 0.867 0.867 0.601
EA. Angle Error 7.416 7.173 8.391 8.267 57.147
Fit Ext. 0.0771 0.0527 0.0791 0.0741 0.1701
Table 4: Comparison with Point2Cyl on DeepCAD dataset.
Table 4 demonstrates the quantitative comparison. Compared with Point2Cyl, the metrics of the extrusion (Ext.) and barrel/based (BB.) segmentation are improved by $5.6\%$ and $3.5\%$, respectively, leading to $13\%$ and $32\%$ reduction on the extrusion axis (EA.) angle and fitting (Fit Ext.) error, respectively. This shows that by supervising the extrusion segmentation with curve loss (type and parameter loss), more geometrical information is introduced, leading to better geometry accuracy. We visualized the sketch predicted by Point2Cyl and Point2Primitive, as shown in Figure 6 (b). It can be seen that because Point2Cyl represents the sketch by SDF, the edges of the sketch profiles are curved, while the counterpart of Point2Primitive is sharp and accurate. Further, we present more visualizations of the predicted sketch by Point2Primitive in the Appendix 2. Figure 6 (c) demonstrates the segmentation results by the two methods. It can be seen that even though the multiple parts of one extrusion are labeled with different extrusion labels (rows 4 and 5), the segmentation network can still learn the geometry representation and cluster the points as the same extrusion. This implies that Point2Primitive can predict simplified primitives that are closer to human design.
5.6 Ablation Study
We perform ablation studies to carefully analyze the DeNoising (DN), the improved transformer decoder, the sketch primitive definition, and the balancing weight in the loss function. All quantitative metrics are measured on the DeepCAD dataset.
No. $\beta$ SkhRep ImpDec $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$
0 2.0 - - 82.19% 72.49% 0.819
1 2.0 - 85.71% 78.61% 0.674
2 1.0 93.16% 84.91% 0.341
3 3.0 93.24% 84.73% 0.337
Ours 2.0 96.14% 86.81% 0.312
Table 5: Ablation study. left panel is the accuracy curves during training; right panel is the quantitative results.
DeNoising. It can be seen from the convergence curve shown in Figure 5 that by employing DeNoising in the training procedure, the ultimate sketch prediction accuracy is improved by $11.4\%$ and $3.1\%$ on $Acc^{SKH}_{t}$ and $Acc^{SKH}_{p}$, respectively. More validation loss curves during training are presented in Appendix 5. From the curves, we can see that by directly predicting the curve parameters and types, the accuracy is extremely high (type error below 0.5% and parameter error below 0.06)
Decoder and Sketch Hierarchy. We replace the improved decoder layer (ImpDec) with the vanilla transformer. Also, we utilize the sketch representation (SkhRep) method in [37], which converts parameter prediction into a classification problem. As shown in Table 5, the $Acc^{SKH}_{t}$ improves by $12.2\%$ by utilizing the ImpDec. Also, the center-prior sketch hierarchy contributes to an $8.4\%$ improvement in $Acc_{p}$. This shows the effectiveness of the proposed improved transformer decoder.
Balancing Weight. Table 5 shows that the balance weight $\beta=2$ performs slightly better than the other conditions, and we set it to our default value.
6 Conclusion
The proposed Point2Primitive can produce CAD reconstruction of high geometric fidelity. The reconstruction is more accurate using curves as the sketch representation instead of SDF. The curve parameters are accurate by direct prediction in an autoregressive way through the improved transformer decoder. In further work, we plan to extend our approach to include more complex modeling commands.
Supplementary Material
1 Evaluation Metrics
1.1 Primitive Type and Parameter Metrics
The command type and command parameter accuracy evaluate how accurate the predicted command is. For each CAD command, we calculate the command type accuracy by:
$ACC_{type}=\frac{1}{N_{c}}\sum\limits_{i=1}^{N_{c}}\Psi[t_{i}=\hat{t}_{i}]$ (7)
where $N_{c}$ denote the total number of the CAD commands, $C_{i}$ and $\hat{C}_{i}$ are the ground truth command type and predicted command type, respectively. $\Psi[i]\in\{0,1\}$ is the indicator function. The command parameter accuracy is calculated by:
$ACC_{param}=\frac{1}{K}\sum\limits_{i=1}^{N_{c}}\sum\limits_{j=1}^{\lvert\hat{p}_{i}\rvert}\Phi[\lvert p_{i,j}-\hat{p}_{i,j}\rvert<\epsilon]\Psi[C_{i}=\hat{C}_{i}]$ (8)
where $K=\sum_{i=1}^{N_{c}}\Psi[C_{i}=\hat{C}_{i}]\lvert p_{i}\rvert$ is the total number of parameters in all correctly recovered commands. $p_{i,j}$ and $\hat{p}_{i,j}$ are ground-truth command parameters and predicted command parameters. $\epsilon$ is the tolerance threshold accounting for the parameter accuracy. In practice, we use $\epsilon=0.01$.
1.2 Extrusion Segmentation IoU
To evaluate the extrusion segmentation, we use the Segmentation (Seg) IoU as the metric. The predicted extrusion segmentation labels are first reordered to correspond with the GT through Hungarian matches. The Seg IoU can be formulated as follows:
$\mathrm{Seg\ IoU}$ $=\frac{1}{K}\sum_{k=1}^{K}RIoU(\mathbbm{1}(\mathbf{\hat{W}}_{:,k}),\mathbf{W}_{:,k})$ (9)
$RIoU(\mathbf{X},\mathbf{Y})$ $=\frac{\mathbf{X}^{\top}\cdot{\mathbf{Y}}}{{\left\|\mathbf{X}\right\|}_{1}+{\left\|\mathbf{Y}\right\|}_{1}-\mathbf{X}^{\top}\cdot{\mathbf{Y}}}$ (10)
where $\mathbf{\hat{W}}_{:,k}$ and $\mathbf{W}_{:,k}$ are the predicted and ground-truth labels of the $k$-th extrusion, respectively. $\mathbbm{1}(\cdot)$ indicates the one-hot conversion.
1.3 Base/Barrel Point Segmentation
The Base/Barrel (BB) point Segmentation accuracy is defined as follows:
$\mathrm{BB\ Acc}=\frac{1}{N}\sum_{i=1}^{N}(\mathbbm{1}(\mathbf{\hat{B}}_{i,:})==\mathbf{B}_{i,:})$ (11)
where $\mathbf{\hat{B}}_{i,:}$ and $\mathbf{B}_{i,:}$ are the predicted and ground-truth barrel label, respectively. $N$ is the number of input point clouds.
1.4 Extrusion Axis Angle Error
Extrusion axis (EA) angle error measures the angle error between the GT and prediction, which is defined as follows:
$\mathrm{EA\ Angle\ Error}=\frac{1}{K}\sum_{k=1}^{K}\arccos(\mathbf{\hat{e}}^{\top}_{k,:}\mathbf{e}_{k,:})$ (12)
where $\mathbf{\hat{e}}^{\top}_{k,:}$ and $\mathbf{e}_{k,:}$ are the predicted and ground-truth extrusion axis, respectively.
1.5 Extrusion Fitting Error
The extrusion fitting (Fit Ext) error measures the average fitting error of each extrusion, which can be defined as follows:
$\mathrm{Fit\ Ext}$ $=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F}$ (13)
$\mathcal{F}$ $\triangleq\frac{1}{N_{\mathbf{P}_{k}}}\sum\left|SDF(\hat{\mathbf{s}}_{k},\hat{\mathbf{S}}_{k})-SDF(\mathbf{s}_{k},\mathbf{S}_{k}))\right|$ (14)
$\hat{\mathbf{S}}_{k}$ $=\prod\limits(\mathbf{P}^{barrel_{k}},\hat{e}_{k,:},\hat{c}_{k,:})$ (15)
where $\hat{c}_{k,:}$ is the predicted extrusion center. ($\prod$($\cdot$)∏ ( ⋅ ) projects per extrusion barrel points using the extrusion axis and center. The inner summation $\mathcal{F}(\mathbf{P},k)$ represents the goodness of the $k$-th extrusion fitting.
1.6 Primitive Number
The primitive number is a metric of reconstructed sequence fidelity. We calculate the $\Delta\#P$ to measure the length difference between the reconstructed and ground-truth sequence, which is defined as follows:
$\Delta\#P=\left|\hat{N}_{c}-N_{c}\right|$ (16)
where $\hat{N}_{c}$ and $N_{c}$ are the predicted and ground-truth primitive numbers.
2 Visualized Sketch Prediction
Figure 7: Visualization of the sketch primitive prediction.
We provide more visualized results of the predicted sketch. Fig. 7 demonstrates some examples of the predicted sketch primitives. It can be seen that the predicted sketches are accurate both in curve type and parameter. However, some parameter errors can be seen in the Figure. Therefore, we develop a post-optimized method to eliminate the parameter errors, which is detailed in Section 5.2.
3 Validation Loss w/o DeNoising
Figure 8: Validation curves with and without DeNoising of the Point2Primitive during training: (a) the class error during training; (b) the L-1 loss of the primitive parameter during training; (c) the distance loss of the primitive parameter during training.
Except for the command type and accuracy curve, Fig. 8 demonstrates the validation loss of the proposed Point2Primitive during training. From the loss curves, we can see that the final loss of the curve type trained with DN is much smaller than the counterpart without DN (0.413 vs 1.1809). As for the curve parameter, the L-1 loss drops by $3.7\%$, and the distance loss drops by $4.9\%$. Also, the convergence time trained with DN is shorter than without DN.
4 The Implementation of the Extrusion Segmentation Network
Figure 9: The implementation of the extrusion segmentation network: (a) the network architecture of the Point2Primitive; (b) the Sketch Point Encoder based on the PointCNN
4.1 Extrusion and Barrel Point Segmentaion
Figure 9 demonstrates the extrusion segmentation network. PointNet++ is used as the encoder and the decoder. The Normal head and Ext head produce the normal prediction $\hat{N}$ and the extrusion segmentation logits $\hat{\mathbf{M}}$. The extrusion segmentation logits $\hat{\mathbf{M}}$ produce the $2K-$classes of each point. The $(2k)$-th elements of each row of the $\hat{\mathbf{M}}$ encode the extrusion label, while the $(2k+1)$-th elements encode the barrel label. Thus the extrusion $\mathbf{\hat{W}}$ and barrel $\mathbf{\hat{B}}$ segmentation logits are formulated as follows:
$\mathbf{\hat{W}}_{:,j}$ $=\mathbf{\hat{M}}_{:,2j}+\mathbf{\hat{M}}_{:,2j+1}$ (17)
$\mathbf{\hat{B}}_{:,0}$ $=\sum_{i}\mathbf{\hat{M}}_{:,2j}$ (18)
$\mathbf{\hat{B}}_{:,1}$ $=\sum_{i}\mathbf{\hat{M}}_{:,2j+1}$ (19)
4.2 Sketch Point Encoder
As shown in Figure 9, the sketch point encoder is developed based on the PointCNN[16]. Instead of using random or fps points, fixed points are set as the representation points. Also, to provide position encodings for the transformer, a learned position encoding is utilized.
5 Post Optimization
5.1 The Batch SDF Computing
We develop a batch SDF computing method to calculate the sketch SDF considering batch input. More specifically, the curves of each sketch will be divided into three curve types ($L,A,C$), and curves of the same type are calculated in batches. We now provide an anonymous GitHub repository to present the SDF computing in https://github.com/AnonymousRepo1234/Point2Primitive/blob/main/SDF_batch_cal.py.
5.2 Curve Fine-tuning
Figure 10: Post optimization of curve prediction.
To produce more accurate curve parameters for CAD reconstruction, we develop a post-optimization method for the predicted curve parameters, as shown in Figure 10. The predicted sketch points can be used to optimize the predicted curve parameters. First, we project the barrel point into sketch points. Then, we calculate the sketch SDF to the sketch points using batch SDF computing. The sketch SDF to its sketch points is set as the loss function. Thirdly, the curve parameters are set as model parameters and updated using the autograd in Pytorch supervised by the SDF. The sketch SDF to its sketch points is supposed to be zero. The algorithm is shown in Algorithm 1.
Algorithm 1 Curve Fine-tuning
$Curves$
$T$ as modified Curves
$Curves_{norm}\leftarrow normalize(Curves)$
$Curves_{norm}\leftarrow sort(Curves_{norm})$triangleright sorted by starting point
$S\leftarrow Curves_{norm}$
$T\leftarrow\emptyset$triangleright as modified Curves
while $S\neq\emptyset$ do
       $x\leftarrow S.front$
       for each curve $l\in S$ do $\triangleright$ modify neighborhood line
             if $\|x.end,l.start\|<Threshold$ then
                    $x.end,l.start\leftarrow mid(x.end,l.start)$
                    $T\leftarrow T\cup\{x\}$
                    $S\leftarrow S\backslash\{x\}$
                    break
             else if $\|x.end,l.end\|<Threshold$ then
                    $x.end,l.end\leftarrow mid(x.end,l.end)$
                    $l\leftarrow reverse(l)$
                    $T\leftarrow T\cup\{x\}$
                    $S\leftarrow S\backslash\{x\}$
                    break
             end if
       end for
end while
$T\leftarrow post\_process(T)$
5.3 Loop Opmitzation
Algorithm 2 Loop Opmitzation
$point\_cloud,Curves$
$params$ as modified sketch curve params
$eval\_points\leftarrow in\_pc(point\_cloud)$triangleright as pc to be formulated
$cmd,params\leftarrow Initialize(sketch)$triangleright cmd as mask of curve type, params as shape of sketch
for each epoch in $epoches$ do
       $line\_loss\leftarrow SDF(eval\_points,params[\newline line\_mask(cmd)])$triangleright calculate line, arc and circle loss
       $arc\_loss\leftarrow SDF(eval\_points,params[\newline arc\_mask(cmd)])$
       $circle\_loss\leftarrow SDF(eval\_points,params[\newline circle\_mask(cmd)])$
       $loss\leftarrow concat(line\_loss,arc\_loss,circle\_loss,-1)$
       $loss.min()$triangleright minimize the last element of loss tuple
       $loss.backward()$
end for
The sketch should contain loops, which are composed of curves connected end to end. To connect the predicted curves into loops, we use an optimization method based on Greedy Algorithm as shown in Algorithm 2.
6 Extrusion Parameters Calculation
We calculate the parameters of the extrusion operation by the obtained barrel points. The parameters of the extrusion operation $E_{i}=(\hat{e_{i}},\hat{t_{i}})$ consists of the extrusion axis $\hat{e_{i}}$, extrusion position $\hat{o_{i}}$, and extrusion extent $\hat{t_{i}}$. Firstly, the optimal extrusion axis of an input extrusion cylinder PC is given by $\hat{e_{i}}=\mbox{argmin}_{e_{i},\|e_{i}\|=1}(e_{i}^{T}H_{\phi}e_{i})$, where:
$H_{\phi}=\textbf{N}^{\top}\Phi_{barr}^{\top}\Phi_{barr}\textbf{N}-\textbf{N}^{\top}\Phi_{base}^{\top}\Phi_{base}\textbf{N}$ (20)
where N denotes the normals of the input points. $\Phi_{barr}=diag(\phi_{barr})$, $\Phi_{base}=diag(\phi_{base})$. $\phi_{base}$ and $\phi_{barr}$ indicate the barrel and base weights (assigned to all points) predicted by the segmentation head, respectively. The solution is given by the eigenvector corresponding to the smallest eigenvalue of $H_{\phi}$.
The extrusion extent $\hat{t}$ is the point projection range on the extrusion axis, which can be calculated by Equation 21.
$\begin{split}&\hat{t}_{min}=\min\limits_{P_{i}\in{P_{barr}}}(e_{i}\cdot(P_{i}-\hat{o}))\\ &\hat{t}_{max}=\max\limits_{P_{i}\in{P_{barr}}}(e_{i}\cdot(P_{i}-\hat{o}))\\ \end{split}$ (21)
7 More Details of the Improved Transformer Decoder
7.1 Command Parameter Noise
For each input training pair $(p_{i},S_{i})$, we add random noises to both their curve types and curve parameters. A random noise $(\Delta x,\Delta y)$ is added to the endpoint coordinates. The values of $(\Delta x,\Delta y)$ are limited between $[0.0,0.2]$ so that the noised endpoint will not shift too far from the ground-truth position. The other primitive parameters $(x_{m_{i}},y_{m_{i}},r_{i})$ are randomly sampled from $(\phi x_{m_{i}},\phi y_{m_{i}},\phi r_{i}),\phi\in[-1,1]$.
For curve type noising, the GT command types are randomly flipped.
7.2 Attention Mask
An attention mask is used to prevent information leakage. There are two reasons for this. Firstly, the matching part might be able to obtain information from the noised curve types and easily infer the target curve types. Secondly, the various noised groups of the same ground-truth curve might exchange information with each other.
Given a sketch contains $N_{c}$ curves. The noised ground-truth curves of all the primitives are first divided into $K$ groups. The denoising part is then denoted as:
$\begin{split}&q=\{g_{0},g_{1},...,g_{K-1}\}\\ &g_{j}=\{q_{0}^{k},q_{1}^{k},...,q_{M-1}^{k}\}\end{split}$ (22)
where $g_{k}$ is the $-k\mbox{-}$th denosing group. $q_{m}^{k}$ is the $-m\mbox{-}$th query in the denoising group. Each denoising group contains $M$ queries where $M$ is the number of commands in one input PC.
The attention mask $A=[a_{ij}]_{W\times{W}}$ can then be formulated as:
$a_{ij}=\left\{\begin{aligned} 1&,\text{ if }j<K\times{M}\text{ and }\lfloor\frac{i}{M}\rfloor\neq\lfloor\frac{j}{M}\rfloor;\\ 1&,\text{ if }j<K\times{M}\text{ and }i\geq K\times{M};\\ 0&,\text{ else }\\ \end{aligned}\right.$ (23)
where $a_{ij}=1$ means the $-i\mbox{-}$th query cannot see the $-j\mbox{-}$th query and $a_{ij}=0$ other wise.
The decoder embeddings are taken as curve embeddings in our model. Therefore, an indicator is appended to the curve embeddings to distinguish the denoising part from the matching part, as shown in Figure 3, where $1$ means the query belongs to the denoising part and 00 means the query belongs to the matching part.
8 Applications
Figure 11: Examples of some practical application: (a) reconstruction of the structural components; (b) reconstruction of the buildings; (c) reconstruction of the small urban agglomeration.
In addition to the public CAD dataset, we also tested our method on some practical datasets. Figure 11 demonstrates some practical applications of the presented CAD reconstruction method.

### v3

@@
 Abstract
-Recovering CAD models from point clouds, especially the sketch-extrusion process, can be seen as the process of rebuilding the topology and extrusion primitives.
-Previous methods utilize implicit fields for sketch representation, leading to shape reconstruction of curved edges.
-In this paper, we proposed a CAD reconstruction network that produces editable CAD models from input point clouds (Point2Primitive) by directly predicting every element of the extrusion primitives.
-Point2Primitive can directly detect and predict sketch curves (type and parameter) from point clouds based on an improved transformer.
-The sketch curve parameters are formulated as position queries and optimized in an autoregressive way, leading to high parameter accuracy.
-The topology is rebuilt by extrusion segmentation, and each extrusion parameter (sketch and extrusion operation) is recovered by combining the predicted curves and the computed extrusion operation.
-Extensive experiments demonstrate that our method is superior in primitive prediction accuracy and CAD reconstruction.
-The reconstructed shapes are of high geometrical fidelity.
-Figure 1: CAD reconstruction pipeline including (a) The Point2Primitive network first segmented the input point cloud into clusters belonging to the same extrusion and corresponding barrel points, and the curves of each extrusion primitive are directly learned and predicted.
-CAD reconstruction can be achieved by combining the reconstructed extrusion sketch and extrusion operations computed from the barrel points, which is easy to edit in downstream tasks.
-(b) The illustrations of the sketch-extrude process, extrusion segmentation, and extrusion primitives.
-The topology describes how extrusion primitives are interrelated.
+Recovering CAD models from point clouds requires reconstructing their topology and sketch-based extrusion primitives.
+A dominant paradigm for representing sketches involves implicit neural representations such as Signed Distance Fields (SDFs).
+However, this indirect approach inherently struggles with precision, leading to unintended curved edges and models that are difficult to edit.
+In this paper, we propose Point2Primitive, a framework that learns to directly predict the explicit, parametric primitives of CAD models.
+Our method treats sketch reconstruction as a set prediction problem, employing a improved transformer-based decoder with explicit position queries to directly detect and predict the fundamental sketch curves (i.e., type and parameter) from the point cloud.
+Instead of approximating a continuous field, we formulate curve parameters as explicit position queries, which are optimized autoregressively to achieve high accuracy.
+The overall topology is rebuilt via extrusion segmentation.
+Extensive experiments demonstrate that this direct prediction paradigm significantly outperforms implicit methods in both primitive accuracy and overall geometric fidelity.
 **footnotetext: Corresponding Author
 1 Introduction
-CAD reconstruction from point clouds is a longstanding objective [10].
-In CAD, reverse engineering encompasses techniques that extract CAD models from existing products by transforming unstructured data (e.g.
-point clouds, meshes, or printed schematics) into structured representations of 2D or 3D geometries.
-The sketch-and-extrude process is one of the most fundamental CAD modeling commands, offering a highly intuitive and versatile way to generate a broad array of shapes, as demonstrated in [36], and investigating the reverse of this process is valuable to the community.
-However, the unordered and unstructured nature of point clouds poses significant challenges for high-level manipulation and efficient editing of their underlying geometries [27].
-Some methods try to solve this problem and reconstruct the shapes through optimization methods [19, 1, 30, 15, 18, 26].
-The recent availability of significant boundary representation (B-Rep) datasets [36, 9, 31] has introduced data-driven frameworks to reverse engineering in the CAD world.
-Numerous neural networks have been developed to segment point clouds into clusters to represent distinct surface regions [13, 32, 11, 21].
-These segmented points are then fitted to surface primitives (either parametric or freeform), allowing the topological assembly of these primitives to reconstruct the CAD models.
-In contrast, we focus on recovering the CAD models that describe the sketch-and-extrude modeling process.
-To this end, we employ a point segmentation network to recover topological structures (extrusion primitives) by segmenting the point clouds into clusters associated with individual extrusion primitives.
-By predicting the curves and extrusion operation of each extrusion primitive, the CAD models can be reconstructed.
-Specifically, the sketch serves as the foundation of parametric CAD modeling, with the quality of the final CAD model depending critically on the precision of its sketches [5].
-In the parametric CAD workflow, close curves constitute the sketch, and 3D models are then constructed by extruding the sketches and grouping these extrusions into structured assemblies.
-Recently, implicit representations of 3D shapes [35, 23], have inspired new approaches that employ SDFs to represent sketches, enabling sketch and CAD reconstruction through SDF encodings [34, 29, 14].
-However, decoding SDF embeddings into sketches typically results in CAD models with unintended curved edges and the reconstructed sketch is not easy to be directly edited.
-Inspired by large language models (LLMs) [4], recent works [37, 39, 22] propose to generate CAD modeling sequences using language-based models.
-However, our findings indicate that LLMs struggle to implicitly learn hierarchical representations of topology (e.g.
-CSG structures) [41] using only reconstruction loss, such as chamfer distance and primitive loss, which can result in limited generalization.
-SkexGen [38] strives to address this challenge by introducing separate sequence encoders for modeling topology and geometry independently to facilitate related design generation.
-In response, we propose representing sketches through their fundamental curves and explicitly recovering their extrusion primitives [41] by decomposing input points into distinct extrusion sets.
-This paper presents a novel approach for reconstructing sketch-and-extrude CAD models using an end-to-end point-to-primitive (Point2Primitive) network.
-Specifically, our method first segments the input point cloud into distinct extrusion primitives and corresponding barrel/base points.
-Each extrusion sketch is then predicted by directly predicting the type and parameter of its curves through an improved transformer network.
-By combining the predicted curves and the extrusion operation calculated from the barrel points, the CAD models can be rebuilt.
-Using curves as sketch representation, our approach reconstructs fundamental shapes in sketches (such as lines, arcs, and circles), resulting in accurate, geometry-preserving, and easily editable 3D CAD models.
-Our primary contributions are summarized as follows:
-  • We propose a CAD reconstruction method based on the Point2Primitive network, which leverages an enhanced transformer to predict sketch and PointNet++ for extrusion segmentation.
-By representing the sketch through its fundamental curves rather than using SDF, our approach achieves higher geometric fidelity.
-  • We introduce an improved transformer decoder along with a sketch hierarchy representation.
-The curve parameters are encoded as position embeddings to guide the attention weight in locating the target.
-The parameters are directly predicted and optimized in an auto-regressive way, leading to high accuracy.
-  • We present a complete data generation pipeline that can produce extrusion segmentation labels, barrel/base segmentation labels, and sketch primitives from the CAD modeling sequence saved in JSON files, which can be used for future research.
-  • We conduct a comprehensive experiment with several CAD reconstruction results to analyze the proposed method thoroughly.
-Also, we reproduced the data generation methods used by other methods for comparison.
-An ablation study is conducted to fully investigate the proposed method.
+Computer-Aided Design (CAD) reconstruction, which aims to convert unstructured data like point clouds into structured, parametric models, is a long-standing goal in computer graphics and engineering (WOS:000072777200021).
+Among various modeling techniques, the sketch-and-extrude process is one of the most fundamental operations, capable of generating a vast array of mechanical parts and industrial designs (willis2021fusion360gallerydataset).
+Therefore, automatically reversing this process, transforming a raw 3D scan into a structured sketch-based CAD representation, holds immense value for digital manufacturing, legacy part recreation, and design modification.
+Despite recent progress driven by large-scale datasets (willis2021fusion360gallerydataset; koch2019abcbigcadmodel), a critical challenge lies in accurately capturing the model’s boundaries.
+The unstructured, noisy nature of point clouds contrasts sharply with the precise, topologically-defined structure of CAD models.
+A dominant paradigm in recent literature has been to represent 2D sketches using implicit neural representations such as Signed Distance Fields (SDFs) (9880393; Ren_Zheng_Cai_Li_Zhang_2022; Li_Guo_Zhang_Yan_2023).
+While these methods are powerful for capturing complex, continuous surfaces, they remain limited in several critical aspects.
+First, SDFs are purely geometric fields that lack semantic structure, offering no direct correspondence to the primitive entities used in CAD.
+Second, as approximation-based representations, they inevitably introduce noise and blur fine geometric details, which makes it difficult to recover sharp corners or well-defined curves.
+Consequently, reconstructed sketches often exhibit smoothed or rounded edges rather than precise boundaries.
+Finally, because SDF-based sketches cannot be directly expressed in terms of editable primitives—lines, circles, and arcs, the lingua franca of CAD software—an additional conversion step is required.
+This transformation not only complicates the workflow but also accumulates parameter errors, further hindering faithful reconstruction and downstream editability.
+The above limitations lead us to a fundamental question:
+Can we bypass the imprecise, intermediate representations and directly predict the explicit, parametric sketch primitives from the point cloud?
+In this work, we affirmatively answer this question by proposing Point2Primitive, a framework grounded in a direct prediction methodology for CAD reconstruction from point clouds.
+Unlike indirect approaches that rely on intermediate surface fitting or voxelization, our method directly targets the underlying parametric sketch primitives.
+Specifically, we observe that many CAD models can be systematically deconstructed into a sequence of extrusion operations, each defined by a two-dimensional sketch—a closed profile formed by geometric primitives such as lines, arcs, and circles—together with an associated three-dimensional operation that specifies the extrusion direction and length.
+Building on this insight, we reformulate the complex reverse engineering problem into two interdependent sub-tasks.
+First, in the stage of topology reconstruction, the input point cloud is partitioned into clusters that correspond to individual extrusions, a process we term extrusion segmentation.
+This step disentangles the overall geometry into meaningful sub-structures, enabling subsequent sketch prediction to operate on localized and semantically coherent point sets.
+Second, in the stage of primitive prediction, we recover the generating two-dimensional sketch for each segmented extrusion body.
+To achieve this, we design an improved transformer-based decoder that refines point features and directly decodes them into parametric primitives.
+The decoder integrates explicit geometric priors by formulating curve parameters as positional queries, which guide the attention mechanism and progressively update predictions across layers.
+This design enhances the accuracy and stability of primitive recovery.
+Point2Primitive could bridge topology recovery and parametric sketch inference within a unified direct-prediction framework.
+By directly predicting the language of CAD, our method produces CAD models that are not only geometrically accurate but also truly editable.
+In summary, our main contributions are summarized as follows:
+  • We introduce Point2Primitive, a new pipeline that directly predicts parametric sketch primitives (i.e., lines, arcs, and circles) from raw point clouds, bypassing implicit fields and preserving sharp, semantically meaningful boundaries for downstream editability.
+  • We decompose reconstruction into (i) extrusion segmentation to recover topology by clustering the point cloud into extrusion bodies, and (ii) primitive prediction with an improved transformer decoder that uses explicit position queries and autoregressive parameter refinement to yield high-precision primitive types and parameters.
+  • Extensive experiments and ablations show consistent gains in boundary fidelity, primitive recovery, and editability over SDF-based and recent learning baselines, demonstrating both accuracy and robustness across diverse CAD geometries.
 2 Related works
@@
 Several learning-based approaches have also been proposed to fit geometric primitives to point clouds.
-These methods first segment the input points into clusters corresponding to the same surface primitives and fit the clustered points to the parametric surface primitives [13, 32, 11, 17] or free surfaces [21].
-Except for surface primitives, UCSG-Net[7] and CSG-stump [28] learn the CSG parse tree that interrelates the predicted geometry primitives to reconstruct shapes.
+These methods first segment the input points into clusters corresponding to the same surface primitives and fit the clustered points to the parametric surface primitives (Li_Sung_Dubrovina_Yi_Guibas_2019; Sharma_Liu_Maji; Le_Sung_Ceylan_Mech_Boubekeur_Mitra_2021; RN21) or free surfaces (RN20).
+Except for surface primitives, UCSG-Net (kania2020ucsgnetunsuperviseddiscovering) and CSG-stump (ren2021csgstumplearningfriendlycsglike) learn the CSG parse tree that interrelates the predicted geometry primitives to reconstruct shapes.
 The reconstructed CAD models are in B-rep format, which is not as easy to edit as the sketch-and-extrude modeling procedure.
@@
 2.2 Sketch-and-extrude CAD Reconstruction
-Some new solutions use SDF to represent sketches following the recent implicit 3D shape representations [23, 35].
-ExtrudeNet [27] and SECAD-Net [14] can learn implicit sketches and differentiable extrusions.
+Some new solutions use SDF to represent sketches following the recent implicit 3D shape representations (Park_Florence_Straub_Newcombe_Lovegrove_2019; wang2022nerf).
+ExtrudeNet (RN93) and SECAD-Net (Li_Guo_Zhang_Yan_2023) can learn implicit sketches and differentiable extrusions.
 Point2Cyl reconstructs 3D shapes through sketch regression supervised by the latent embeddings of its SDF.
@@
 Further transformation from SDF to sketch further brings in extra parameter errors.
-In this paper, we propose to represent a sketch using its curves and directly learn and predict the curve types and parameters from the point clouds in an object detection manner [33].
-To this end, we propose the center-prior primitive definition (detailed in Section 3) and the improved transformer decoder (detailed in Section 4.4) to achieve precise and editable sketch reconstruction, which makes the final CAD reconstruction easy to be edited.
-Also, the most related work to us is the Point2Cyl, and we conduct an extra comparison as briefed in Section .
 2.3 CAD Generation
-Following Large Language Model (LLM) [4], [37] is the first to explore the CAD sequence generation of 3D shapes based on VAE [8].
-[39] utilizes a vector quantized VAE to generate modeling sequences.
-[22] propose to generate CAD models step-by-step based on multimodal diffusion.
-It should be noted that the CSG structure [41] is broken when tokenizing the CAD modeling sequence.
-We find that the implicit representation learning of the topology is challenging for LLM if only supervised by geometry loss (such as Chamfer Distance) or command accuracy.
-SkexGen [38] propose to use two sequence encoders separately for topology and geometry learning.
-However, CAD generation methods leveraging VAE [8] structure and language model decoder still limit its performance on reconstruction tasks because of poor generalization.
-Also, language models struggle with numerical representation and mathematics [40], causing inaccuracies in command parameters and limited performance beyond the training dataset (detailed in Section 5.4).
-In this paper, we present a CAD reconstruction method based on the Point2Primitive network.
-We explicitly recover the topology through point segmentation and directly predict the sketch curves and extrusion operation.
-By explicit precise primitive parameter prediction, the CAD reconstructions are of high geometry fidelity and good generalization.
-3 Preliminaries
-In CAD terminology, the sketch consists of multiple closed curves $S_{i}$.
-There are three types of curves: line $\mathit{L}$, arc $\mathit{A}$, and circle $\mathit{R}$ in our sketch hierarchy.
-Therefore, the sketch $S_{i}=[c_{1}^{i},...,c_{N}^{i}]$, where $c_{j}^{i}\in(\mathit{L},\mathit{A},\mathit{R})$ for all $j\in[1,...,N]$, can be seen as a set of curves.
-The goal of predicting sketch is to train the parameters $\theta$ of the Point2Primitive network $f_{\theta}$ on the training pair $({p_{i}},\hat{S}_{i})$ to make a good approximator of the set prediction as follows:
-$f_{\theta}\left(\{p_{k}\mid p_{k}\in P_{\mathbb{E}_{i}}\}\right)\approx\left\{c_{1}^{i},...,c_{j}^{i}\mid c_{j}^{i}\in(\mathit{L},\mathit{A},\mathit{C})\right\}$ (1)
-where $P_{\mathbb{E}_{i}}\in R^{N_{E_{i}}\times{3}}$ is the points of the $i\mbox{-}th$ extrusion $\mathbb{E}_{i}$.
-Curve Type Parameters
-$L$ $(Line)$ $x_{m}$ $y_{m}$ $x_{1}$ $y_{1}$1 -1
-$A$ $(Arc)$ $x_{m}$ $y_{m}$ $x_{1}$ $y_{1}$ $x_{2}$ $y_{2}$
-$R$ $(Circle)$ $x_{c}$ $y_{c}$ $r$1 -1 -1
-Table 1: Curve definition. We provide the types and vector representations of curves used in this work. Unused parameters in curve vector are padded with $-1$.
-To work with the improved transformer decoder, we formulate the definition of curve parameter in a centra-prior way, as shown in Table 1. Each primitive parameter is primarily presented by its center.
-Line is represented with its midpoint ($x_{m},y_{m}$) and another endpoint ($x_{1},y_{1}$).
-Arc is defined with its midpoint ($x_{m},y_{m}$), start points ($x_{1},y_{1}$) and endpoints ($x_{2},y_{2}$).
-The circle parameter consists of the center point ($x_{c},y_{c}$) and radius $r$.
-Each curve $C_{i}=(t_{i},\rho_{i})$ is defined by its primitive types $t_{i}\in{\mathbb{R}}$ and primitive parameters $\rho_{i}\in\mathbb{R}^{1\times{6}}$ with unused parameters are padded with $-1$.
-Therefore, the first two elements of the parameters $\rho_{i}$ can be encoded as position embeddings to guide the attention in the improved transformer decoder as briefed in Section 4.4
-In the sketch-extrude process, a sketch $S_{i}$ is extruded by an operation $E_{i}$ to form an extrusion $\mathbb{E}_{i}=\left[S_{i},E_{i}\right]$, and multiple such extrusions collectively form a solid shape $\mathbb{S}=\{\mathbb{E}_{1},...,\mathbb{E}_{K}\}$.
-The shape $\mathbb{S}$, therefore, can be represented by the SDF defined by the collection of the extrusions.
-In our method, the $\{\mathbb{E}_{1},...,\mathbb{E}_{i}\}$ is reconstructed via point segmentation, while the sketch $S_{i}$ is predicted through an improved transformer network.
-4 Point2Primtive
-4.1 Overview
-We propose Point2Primitive, a deep model for CAD reconstruction from point clouds.
-As illustrated in Figure 2, our method utilizes a point segmentation network to cluster points that belong to the same extrusion.
-Base/barrel points are also classified during extrusion segmentation.
-Then an improved transformer is introduced to predict the curve type and parameters of each extrusion.
-As defined in Equation 1, curve prediction can be seen as a set prediction problem.
-Therefore, we use the transformer network to solve the set-to-set problem and make some improvements to the transformer decoder, as shown in Figure 2.
-The transformer encoder is used to refine the barrel point features from the segmentation (Seg) head.
-The Seg head is detailed in Appendix 4.2, where the barrel point encoder is developed based on the PointCNN [16] with fixed grid points as representation points instead of random points.
-4.2 CAD Dataset Generation
-In this paper, we propose to generate the training dataset based on the Signed Distance Field (SDF).
-An SDF $SDF(p_{eval},b)$ is a continuous function that, for a given spatial eval point $p_{eval_{i}}$, outputs the eval point’s distance to the closest boundary defined by $b$, whose sign denotes whether the point is inside (negative) or outside (positive) of the watertight surface.
-The data generation pipeline can convert any CAD sequence saved in JSON files following the style of the Fusion360 Gallery dataset [36] to the training data pairs.
-Given an extrusion point set $P_{\mathbb{E}_{i}}\in R^{N_{\mathbb{E}_{i}}\times{3}}$ and its sketch $S_{i}$.
-We first project the points $P_{\mathbb{E}_{i}}$ to the sketch plane as eval points according to the transform matrix $W_{i}$ defined by its extrusion axis $e_{i}$.
-Then, its SDF is calculated to the boundary defined by the sketch $S_{i}$.
-The batch SDF calculation will be detailed in the Appendix 5.1.
-The barrel label $b_{i}$ of each point can be defined as follows:
-$b_{i}=\text{bool}(\text{SDF}(W_{i}\cdot{P_{\mathbb{E}_{i}}},S_{i})\leq\text{thresh})$ (2)
-where $thresh$ controls the label precision and $\text{thresh}=0.01$ in this paper.
-As for the Extrusion Label, each extrusion mesh is first built using pythonOCC [24] by the CAD sequence.
-Then, the SDF of the input points to each extrusion boundary defined by its mesh is calculated.
-The extrusion label of each point is set as its closest extrusion mesh.
-4.3 Extrusion Segmentation
-Our method utilizes a point segmentation network to cluster points and rebuild the topology, as shown in Figure 2.
-The Point2Primitive is insensitive to the implementation of the point segmentation network.
-We choose the PointNet++ [25] as the point encoder and the decoder, which is detailed in Appendix 4.1.
+Early attempts explored CAD sequence generation with VAE-based models (Wu_Xiao_Zheng_2021; kingma2022autoencodingvariationalbayes) and their extensions such as vector-quantized VAE (xu2023hierarchicalneuralcodingcontrollable) or multimodal diffusion frameworks (MaCLLZ24).
+However, tokenizing CAD sequences often disrupts the underlying CSG structure (yu2021caprinetlearningcompactcad), making it difficult for language models to capture topology.
+While SkexGen (Xu_Willis_Lambourne_Cheng_Jayaraman_Furukawa_2022) addressed this by separately encoding topology and geometry, existing VAE–LM hybrids still generalize poorly to reconstruction tasks.
+Moreover, LLMs struggle with numerical precision (yang2024llmi3dempoweringllm3d), leading to parameter inaccuracies and limited transferability beyond training data.
+3 Method
+3.1 Preliminaries
+A 2D CAD sketch can be modeled as a finite sequence of geometric primitives $S_{i}=(\,c^{(i)}_{1},\ldots,c^{(i)}_{N_{i}}\,)$, where each primitive $c^{(i)}_{j}=(t^{(i)}_{j},\rho^{(i)}_{j})$ comprises a type $t^{(i)}_{j}\in\mathcal{T}$ and a parameter vector $\rho^{(i)}_{j}\in\mathbb{R}^{6}$.
+We adopt three primitive types $\mathcal{T}=\{\mathsf{L},\mathsf{A},\mathsf{C}\}$ for line, arc, and circle, respectively.
+Given the point set of the $i$-th extrusion, $\textbf{P}_{\mathcal{E}_{i}}\in\mathbb{R}^{N_{\mathcal{E}_{i}}\times 3}$, the network $f_{\theta}$ is trained to predict the sketch as a set/sequence of primitives:
+$f_{\theta}\!\left(\{\,\mathbf{p}_{k}\mid\mathbf{p}_{k}\!\in\!\textbf{P}_{\mathcal{E}_{i}}\,\}\right)\;\approx\;S_{i}\;=\;(\,c^{(i)}_{1},\ldots,c^{(i)}_{N_{i}}\,),\quad c^{(i)}_{j}=(t^{(i)}_{j},\rho^{(i)}_{j}),\;t^{(i)}_{j}\!\in\!\{\mathsf{L},\mathsf{A},\mathsf{C}\}.$ (1)
+Center-prior parameterization.
+We use a center-prior parameterization with a fixed 6-D vector $\rho=(\rho_{1},\ldots,\rho_{6})$ and pad unused entries by 1-1.
+Line ($\mathsf{L}$): midpoint $\mathbf{m}=(x_{m},y_{m})$ and one endpoint $\mathbf{d}_{1}=(x_{1},y_{1})$, encoded as $\rho=(x_{m},y_{m},x_{1},y_{1},-1,-1)$.
+Arc ($\mathsf{A}$): midpoint $\mathbf{m}=(x_{m},y_{m})$, start point $\mathbf{s}=(x_{1},y_{1})$, and end point $\mathbf{d}_{2}=(x_{2},y_{2})$, encoded as $\rho=(x_{m},y_{m},x_{1},y_{1},x_{2},y_{2})$.
+Circle ($\mathsf{C}$): center $\mathbf{c}=(x_{c},y_{c})$ and radius $r$, encoded as $\rho=(x_{c},y_{c},r,-1,-1,-1)$.
+Unless otherwise stated, we refer to a primitive by $c=(t,\rho)$.
+The first two entries of $\rho$ (i.e., the center or midpoint) are further used as positional embeddings to guide attention in the improved transformer decoder (see Sec.
+3.4).
+Sketch–extrude–shape relation.
+An extrusion is denoted by $\mathcal{E}_{i}=[S_{i},E_{i}]$, where $E_{i}$ is the associated operation.
+A solid shape is a finite collection of extrusions, $\mathbb{S}=\{\mathcal{E}_{1},\ldots,\mathcal{E}_{K}\}$, and is represented by the signed distance field (SDF) induced by these extrusions.
+In our pipeline, $\{\mathcal{E}_{1},\ldots,\mathcal{E}_{i}\}$ is recovered via point segmentation, while the current sketch $S_{i}$ is predicted with the improved transformer decoder.
 Figure 1: Overview of Point2Primitive.
@@
 In sketch prediction, the point features from the Seg head are first refined using a vanilla transformer encoder and decoded into curves through the improved transformer decoder.
+Based on the above preliminaries, we present Point2Primitive, a direct sketch primitives prediction method for CAD reconstruction from point clouds.
+As shown in Figure 1, our method first segments points into extrusions, classifies base/barrel points, and then applies an improved transformer to predict curve types and parameters as a set-to-set problem.
+The segmentation head builds on PointCNN with fixed grid points, while the transformer refines features and enhances prediction accuracy.
+3.2 CAD Dataset Generation
+In this paper, we propose to generate the training dataset based on the Signed Distance Field (SDF).
+An SDF $\texttt{SDF}(p_{\text{eval}},b)$ is a continuous function that, for a given spatial eval point $p_{\text{eval}_{i}}$, outputs the eval point’s distance to the closest boundary defined by $b$, whose sign denotes whether the point is inside (negative) or outside (positive) of the watertight surface.
+The data generation pipeline can convert any CAD sequence saved in JSON files following the style of the Fusion360 Gallery dataset (willis2021fusion360gallerydataset) to the training data pairs.
+Given an extrusion point set $\textbf{P}_{\mathcal{E}_{i}}\in R^{N_{\mathcal{E}_{i}}\times{3}}$ and its sketch $S_{i}$.
+We first project the points $\textbf{P}_{\mathcal{E}_{i}}$ to the sketch plane as eval points according to the transform matrix $W_{i}$ defined by its extrusion axis $e_{i}$.
+Then, its SDF is calculated to the boundary defined by the sketch $S_{i}$.
+The batch SDF calculation will be detailed in the Appendix F.1.
+The barrel label $b_{i}$ of each point can be defined as follows:
+$b_{i}=\texttt{Bool}(\texttt{SDF}(W_{i}\cdot{\textbf{P}_{\mathcal{E}_{i}}},S_{i})\leq{\rm thresh})\,$ (2)
+where $\rm thresh$} controls the label precision and ${\rm thresh}=0.01$ in this paper.
+As for the Extrusion Label, each extrusion mesh is first built using pythonOCC (thomas_paviot_2022_7471333) by the CAD sequence.
+Then, the SDF of the input points to each extrusion boundary defined by its mesh is calculated.
+The extrusion label of each point is set as its closest extrusion mesh.
+3.3 Extrusion Segmentation
+To decompose the raw point cloud into semantically meaningful extrusion bodies, we design an extrusion segmentation network that clusters points belonging to the same extrusion.
+As illustrated in Figure 1, the segmentation module is implemented using a PointNet++ encoder–decoder backbone (Qi_Yi_Su_Guibas_2017), followed by two prediction heads.
+Specifically, the encoder extracts hierarchical features from the input point cloud, while the decoder progressively recovers point-wise embeddings.
+On top of these embeddings, two parallel branches are employed: the Normal Head predicts point-wise surface normals, and the Extrusion Segmentation Head predicts extrusion-related classes.
+The Ext Head outputs logits $\hat{\mathbf{M}}$ corresponding to $2K$ categories, where each extrusion is represented by two complementary channels.
+For the $j$-th extrusion, the $(2j)$-th channel encodes the extrusion body assignment, whereas the $(2j+1)$-th channel encodes its associated barrel points.
+In this way, extrusion logits $\hat{\mathbf{W}}$ (for generating the extrusion $\textbf{p}_{k}$) and barrel logits $\hat{\mathbf{B}}$ (for defining the the boundary $b$) can be computed as:
+$\hat{\mathbf{W}}_{:,j}$ $=\hat{\mathbf{M}}_{:,2j}+\hat{\mathbf{M}}_{:,2j+1},$ (3)
+$\hat{\mathbf{B}}_{:,0}$ $=\sum_{j}\hat{\mathbf{M}}_{:,2j},\quad\text{and}\quad\hat{\mathbf{B}}_{:,1}=\sum_{j}\hat{\mathbf{M}}_{:,2j+1}.$ (4)
+Through this design, the network jointly captures extrusion topology and barrel boundary information, which facilitates accurate decomposition of the complex shape into extrusion units.
+Please refer to Appendix E for more implementation detail.
 3.4 Improved Transformer Decoder
@@
 Parameter as Position Encoding.
-The curve parameters are formulated in a center-prior format detailed in Section 3, so the curve parameters can be formulated as position encoding $PE_{i}$ to guide the attention.
+The curve parameters are formulated in a center-prior format detailed in Section 3.1, so the curve parameters can be formulated as position encoding $\textbf{PE}_{i}$ to guide the attention.
 Given the curve parameters $\rho^{i}_{j}$ as the parameters of the $j\mbox{-}th$ curve in the $i\mbox{-}th$ sketch $S_{i}$.
-The corresponding positional queries $PE_{j}^{i}$ are generated by:
-$PE_{j}^{i}=\text{MLP}(f_{pe}(\rho^{i}_{j}))=\text{MLP}(f_{pe}(x_{m_{j}}^{i},y_{m_{j}}^{i}))$ (3)
+The corresponding positional queries $\textbf{PE}_{j}^{i}$ are generated by:
+$\textbf{PE}_{j}^{i}=\texttt{MLP}(f_{\text{pe}}(\rho^{i}_{j}))=\texttt{MLP}(f_{\text{pe}}(x_{m_{j}}^{i},y_{m_{j}}^{i})),$ (5)
 where $f_{pe}$ is the function that generates the sinusoidal embeddings from the curve center.
-Multi-layer perception (MLP) and ReLU activations are used to produce the position encodings $PE_{j}^{i}$.
+Multi-layer perception (MLP) and ReLU activations are used to produce the position encodings $\textbf{PE}_{j}^{i}$.
 Direct Curve Parameter Prediction.
@@
 The layer-by-layer update is based on the decoder output as follows:
-$\delta x_{m_{j}}^{i},\delta y_{m_{j}}^{i},\delta p_{2_{j}}^{i},\delta p_{3_{j}}^{i},\delta p_{4_{j}}^{i},\delta p_{5_{j}}^{i}=\text{MLP}(O_{n})$ (4)
+$\delta x_{m_{j}}^{i},\delta y_{m_{j}}^{i},\delta p_{2_{j}}^{i},\delta p_{3_{j}}^{i},\delta p_{4_{j}}^{i},\delta p_{5_{j}}^{i}=\texttt{MLP}(O_{n}),$ (6)
 where $O_{n}$ is the output of the $n\mbox{-}th$ decoder layer.
-The curve parameters are updated to $[x_{m_{j}}^{i^{\prime}},y_{m_{j}}^{i^{\prime}},p_{2_{j}}^{i^{\prime}},p_{3_{j}}^{i^{\prime}},p_{4_{j}}^{i^{\prime}},p_{5_{j}}^{i^{\prime}}]$ in each layer, as shown in Figure 3.
+The curve parameters are updated to $x_{m_{j}}^{i^{\prime}},y_{m_{j}}^{i^{\prime}},p_{2_{j}}^{i^{\prime}},p_{3_{j}}^{i^{\prime}},p_{4_{j}}^{i^{\prime}},p_{5_{j}}^{i^{\prime}}]$ in each layer, as shown in Figure 2.
 All intermediate parameter outputs are supervised by ground truth.
-DeNoising [12] is utilized to accelerate the convergence.
-4.5 Loss Function
-The Point2Primitive is trained using a multi-task, non-convex objective composed of extrusion segmentation loss $\mathcal{L}_{\mathbb{E}}$, base-barrel classification loss $\mathcal{L}_{{bb}}$, normal $\mathcal{L}_{norm}$ and sketch $\mathcal{L}_{skh}$ prediction losses:
-$\mathcal{L}=\mathcal{L}_{\mathbb{E}}+\mathcal{L}_{{bb}}+\mathcal{L}_{norm}+\mathcal{L}_{skh}$ (5)
-where the cross entropy loss is used in $\mathcal{L}_{\mathbb{E}}$ and $\mathcal{L}_{{bb}}$ while L-1 loss is used in $\mathcal{L}_{norm}$.
-The sketch prediction loss $\mathcal{L}_{skh}$ can be formulated as:
-$\mathcal{L}_{skh}$ $=\sum\limits_{i=1}^{N_{C}}l_{focal}(\hat{t}_{i},t_{i})+\beta\sum\limits_{i=1}^{N_{C}}\sum\limits_{j=1}^{N_{p}}\mathcal{L}_{p}(\hat{\rho}_{i,j},\rho_{i,j})$ (6)
-$\mathcal{L}_{p}$ $=L_{1}(\hat{\rho}_{i,j},\rho_{i,j})+L_{2}(\hat{\rho}_{i,j},\rho_{i,j})$
-where $l_{focal}(\cdot,\cdot)$ represents the focal loss [20] and $L_{1}(\cdot,\cdot)$ and $L_{2}(\cdot,\cdot)$ represents the L1 and L2 loss, respectively.
+We further utilize query denoising (Li_Zhang_Liu_Guo_Ni_Zhang) to accelerate the convergence.
+3.5 Training Objective
+The Point2Primitive  is trained using a multi-task objective composed of extrusion segmentation loss $\mathcal{L}_{\text{ext}}$, base-barrel classification loss $\mathcal{L}_{\text{bb}}$, normal $\mathcal{L}_{\text{norm}}$ and sketch $\mathcal{L}_{\text{skh}}$ prediction losses:
+$\mathcal{L}=\mathcal{L}_{\text{ext}}+\mathcal{L}_{\text{bb}}+\mathcal{L}_{\text{norm}}+\mathcal{L}_{\text{skh}},$ (7)
+where the cross entropy loss is used in $\mathcal{L}_{\text{ext}}$ and $\mathcal{L}_{\text{bb}}$ while L-1 loss is used in $\mathcal{L}_{\text{norm}}$.
+The sketch prediction loss $\mathcal{L}_{\text{skh}}$ can be formulated as:
+$\mathcal{L}_{skh}$ $=\sum\limits_{i=1}^{N_{C}}\mathcal{L}_{\text{focal}}(\hat{t}_{i},t_{i})+\beta\sum\limits_{i=1}^{N_{C}}\sum\limits_{j=1}^{N_{p}}\mathcal{L}_{\text{p}}(\hat{\rho}_{i,j},\rho_{i,j}),$ (8)
+$\mathcal{L}_{p}$ $=\mathcal{L}_{1}(\hat{\rho}_{i,j},\rho_{i,j})+\mathcal{L}_{1}(\hat{\rho}_{i,j},\rho_{i,j}),$
+where $\mathcal{L}_{\text{focal}}(\cdot,\cdot)$ represents the focal loss (Lin_Goyal_Girshick_He_Dollar_2017) and $\mathcal{L}_{1}(\cdot,\cdot)$ and $\mathcal{L}_{2}(\cdot,\cdot)$ represents the L1 loss and L2 loss, respectively.
 Both denoising loss and curve Loss use the sketch prediction loss $\mathcal{L}_{skh}$, as shown in Figure 1.
@@
 4.1 Experimental Setups
-Datasets. To evaluate the proposed model, we conduct experiments on DeepCAD [37] and Fusion 360 Gallery [36]. The training data is generated as described in Sec. 4.2.
+Datasets.
+To evaluate the proposed model, we conduct experiments on DeepCAD (Wu_Xiao_Zheng_2021) and Fusion 360 Gallery (willis2021fusion360gallerydataset).
+The training data is generated as described in Sec.
+3.2.
 Implementation Details.
-We use PointNet++[25] for extrusion segmentation.
-The Point2Primitive is trained for 400 epochs with a total batch size of 64 and learning rate (lr) $51e-$ 5 with linear warmup and step reduce lr scheduler.
-As for the DeNoising settings, the noise rate for the curve type and parameters is $0.2$ and $0.3$, respectively, with 5 DeNoising groups.
+We use PointNet++ (Qi_Yi_Su_Guibas_2017) for extrusion segmentation.
+The Point2Primitive  is trained for 400 epochs with a total batch size of 64 and learning rate (lr) 1e51e-5 with linear warmup and step reduce lr scheduler.
+As for the query denoising settings, the noise rate for the curve type and parameters is $0.2$ and $0.3$, respectively, with 5 denoising groups.
 The transformer encoder and decoder contain 6 and 8 transformer layers, respectively, with eight attention heads and latent model dimension 256.
 Evaluation Metrics.
-We adopt command type accuracy ($Acc^{SKH}_{t}$) and parameter accuracy ($Acc^{SKH}_{p}$) for quantitative evaluations of the sketch prediction following [5].
-For networks that represent sketch by the implicit nueral field, the tools for convert sketch SDF to command are developed based on the tool by SECAD [14].
-Furthermore, following [3], we also report chamfer distance ($CD$), edge chamfer distance ($ECD$), normal consistency ($NC$), and the number of generated primitives ($\Delta\#P$) to measure the quality of the recovered 3D geometry, more details of the quantitative metrics are shown in the Appendix 1.
+We adopt command type accuracy ($Acc^{SKH}_{t}$) and parameter accuracy ($Acc^{SKH}_{p}$) for quantitative evaluations of the sketch prediction following (GONZALEZLLUCH2021103460).
+For networks that represent sketch by the implicit nueral field, the tools for convert sketch SDF to command are developed based on the tool by SECAD (Li_Guo_Zhang_Yan_2023).
+Furthermore, following Chen_Zhang_2019, we also report chamfer distance ($CD$), edge chamfer distance ($ECD$), normal consistency ($NC$), and the number of generated primitives ($\Delta\#P$) to measure the quality of the recovered 3D geometry, more details of the quantitative metrics are shown in the Appendix B.
 4.2 Main Results
-[b] DeepCAD [37] Fusion 360 Gallery [36] Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$ IR #$\#\Delta P$ $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$ IR #$\#\Delta P$ Methods of Representing Sketch with Implicit SDF Encoding ExtrudeNet [29] 34.58% 31.71% 0.379 0.962 0.849 24.19% 34.34 37.81% 34.39% 0.671 0.808 0.809 23.83% 27.47 Point2Cyl [34] 41.37% 39.41% 0.489 1.027 0.819 3.91% 26.18 42.98% 41.16% 0.529 0.769 0.769 3.87% 26.96 SECAD-Net [14] 45.91% 41.37% 0.341 0.868 0.861 8.03% 32.78 46.82% 42.97% 0.449 0.684 0.813 7.82% 28.61 Generation Methods Based on Language Model DeepCAD [37] 82.61% 73.36% 0.898 1.883 0.823 14.12% 5.89 77.31% 69.81% 7.128 8.729 0.719 13.18% 7.47 HNC-CAD [39] 84.31% 76.71% 0.827 1.064 0.846 6.01% 4.43 80.62% 72.19% 4.381 5.571 0.748 5.92% 6.14 Methods Based on Primitive Fitting UCSG-Net [7] - - 1.849 1.174 0.820 - 14.19 - - 0.952 1.277 0.770 - 11.17 CSG-Stump [28] - - 3.031 0.755 0.828 - 19.46 - - 0.781 0.991 0.744 - 13.08 Our Method 96.14% 86.81% 0.312 0.581 0.897 3.71% 4.14 94.17% 83.52% 0.392 0.571 0.839 3.62% 5.17
+[b] DeepCAD Fusion 360 Gallery Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD ECDECD NCNC IR $\Delta P$ $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD ECDECD NCNC IR $\Delta P$ Methods of Representing Sketch with Implicit SDF Encoding ExtrudeNet 34.58% 31.71% 0.379 0.962 0.849 24.19% 34.34 37.81% 34.39% 0.671 0.808 0.809 23.83% 27.47 Point2Cyl 41.37% 39.41% 0.489 1.027 0.819 3.91% 26.18 42.98% 41.16% 0.529 0.769 0.769 3.87% 26.96 SECAD-Net 45.91% 41.37% 0.341 0.868 0.861 8.03% 32.78 46.82% 42.97% 0.449 0.684 0.813 7.82% 28.61 Generation Methods Based on Language Model DeepCAD 82.61% 73.36% 0.898 1.883 0.823 14.12% 5.89 77.31% 69.81% 7.128 8.729 0.719 13.18% 7.47 HNC-CAD 84.31% 76.71% 0.827 1.064 0.846 6.01% 4.43 80.62% 72.19% 4.381 5.571 0.748 5.92% 6.14 Methods Based on Primitive Fitting UCSG-Net - - 1.849 1.174 0.820 - 14.19 - - 0.952 1.277 0.770 - 11.17 CSG-Stump - - 3.031 0.755 0.828 - 19.46 - - 0.781 0.991 0.744 - 13.08 Ours 96.14% 86.81% 0.312 0.581 0.897 3.71% 4.14 94.17% 83.52% 0.392 0.571 0.839 3.62% 5.17
 Table 1: Evaluation results on the test set of the CAD reconstruction dataset. CAD reconstruction and generation methods are selected for comparison.
-To demonstrate its effectiveness, we compare the proposed model to multiple existing methods, covering the methods of representing sketch with SDF encoding [14, 29, 34], the generation methods based on language model [37, 39], and methods based on primitive fitting [7, 28].
+To demonstrate its effectiveness, we compare the proposed model to multiple existing methods, covering the methods of representing sketch with SDF encoding (Li_Guo_Zhang_Yan_2023; Ren_Zheng_Cai_Li_Zhang_2022; 9880393), the generation methods based on language model (Wu_Xiao_Zheng_2021; xu2023hierarchicalneuralcodingcontrollable), and methods based on primitive fitting (kania2020ucsgnetunsuperviseddiscovering; ren2021csgstumplearningfriendlycsglike).
 The direct reconstruction results are visualized for comparison.
@@
 This shows that by directly recovering curves of the extrusion primitive, our method can achieve better performance than the other extrusion-segmentation methods that represent the sketch with an implicit field.
-As for the number of generated primitives, our method is closer to human designs (The smallest #$\#\Delta P$ value).
+As for the number of generated primitives, our method is closer to human designs (The smallest $\Delta P$ value).
 This improvement is due to the explicit fine-grained reconstruction from each extrusion primitive; every parameter of the curves and extrusion operations are all predicted and optimized.
 4.3 Visualization Results
-Figure 4: Visual comparison between reconstruction results on the DeepCAD (left) and Fusion360 gallery dataset (right).
+Figure 3: Visual comparison between reconstruction results on the DeepCAD (the first 4 lines) and Fusion360 gallery dataset (the last 4 lines).
+Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD ECDECD NCNC $\Delta P$
+Methods of Representing Sketch with Implicit SDF Encoding
+ExtrudeNet 28.39% 25.61% 0.614 1.117 0.776 36.14
+Point2Cyl 34.47% 30.25% 0.518 1.065 0.791 27.96
+SECAD-Net 36.61% 31.43% 0.437 1.079 0.806 34.18
+Generation Method Based on Language Model
+DeepCAD 61.41% 43.71% 5.919 6.883 0.708 7.16
+HNC-CAD 63.39% 53.29% 6.864 7.064 0.711 6.56
+Methods Based on Primitive Fitting
+UCSG-Net - - 2.146 1.273 0.797 14.81
+CSG-Stump - - 3.681 0.958 0.804 20.16
+Ours 94.67% 83.81% 0.410 0.607 0.819 4.97
+Table 2: Evaluation results on the Augmented DeepCAD dataset.
 Figure 3 demonstrates the visualized reconstruction results on the DeepCAD and Fusion 360 Gallery datasets, respectively.
@@
 Still, the edges are critically more accurate than the ExtrudeNet, whose reconstructions are less complete.
+Please refer to Appendix I for more visualizations.
 4.4 Comparison on Augmented DeepCAD Dataset
@@
 We train each method on the original DeepCAD dataset and test all the methods on the augmented DeepCAD test sets.
-Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$$\#\Delta P$
-Methods of Representing Sketch with Implicit SDF Encoding
-ExtrudeNet [29] 28.39% 25.61% 0.614 1.117 0.776 36.14
-Point2Cyl [34] 34.47% 30.25% 0.518 1.065 0.791 27.96
-SECAD-Net [14] 36.61% 31.43% 0.437 1.079 0.806 34.18
-Generation Method Based on Language Model
-DeepCAD [37] 61.41% 43.71% 5.919 6.883 0.708 7.16
-HNC-CAD [39] 63.39% 53.29% 6.864 7.064 0.711 6.56
-Methods Based on Primitive Fitting
-UCSG-Net [7] - - 2.146 1.273 0.797 14.81
-CSG-Stump [28] - - 3.681 0.958 0.804 20.16
-Our Method 94.67% 83.81% 0.410 0.607 0.819 4.97
-Table 3: Evaluation results on the Augmented DeepCAD dataset
 Figure 4: Visual comparison between reconstruction results on the augmented DeepCAD dataset.
+Metrics Ours Point2Cyl
+w/o $\mathcal{L}_{\text{skh}}$ w/ w/o $\mathcal{L}_{\text{skh}}$ w/
+Ext. Seg IoU 0.814 0.862 0.801 0.817
+BB. Seg Acc 0.902 0.902 0.867 0.867
+EA. Angle Err $\downarrow$ 7.416 7.173 8.391 8.267
+Fit Ext. $\downarrow$ 0.0771 0.0527 0.0791 0.0741
+(a) Quantitative comparison on DeepCAD.
+(b) There constructed sketches by Point2Cyl and Point2Primitive.
+(c) The extrusion segmentation by Point2Cyl and Point2Primitive.
+Figure 5: Comparison with Point2Cyl.
 Table 2 demonstrates the quantitative results of the augmented DeepCAD dataset.
@@
 Some reconstruction results by the language model preserve similarities to the original shape, leading to false reconstruction.
+Please refer to Appendix I for more visualizations.
 The generated results are similar to the ground truth but with low geometry accuracies.
@@
 4.5 Detailed Comparison with Point2Cyl
-Figure 6: Comparision with the Point2Cyl: (a) The network architecture of Point2Cyl (left) and our Point2Primitive (right); (b) The reconstructed sketch by Point2Primitive and Point2Cyl; (c) The extrusion segmentation by Point2Primitive and Point2Cyl.
-Point2Cyl is the most related method to the proposed Point2Primitive, so we conduct extra comparisons with Point2Cyl.
-Figure 6 demonstrates the architecture difference between Point2Cyl and the proposed Point2Primitive.
-Instead of using IGR [6] and PointNet[2] to encode SDF for sketch SDF prediction, an improved transformer is utilized in the Point2Primtive to convert barrel points to curves directly.
-Thus, the sketch loss is the curve prediction loss rather than the regression loss of the sketch SDF encodings, which introduces more explicit geometry information.
-Also, we utilize the technique based on the jet fitting [1] (for sketch prediction) and Hough Transform [19] (for extrusion axis) as the optimization methods for comparison.
-Metrics Ours Ours Point2Cyl[34] Point2Cyl[34] JF[1]
-w/o $\mathcal{L}_{sketch}$ w/o $\mathcal{L}_{sketch}$ HT[19]
-Ext. Seg IoU 0.814 0.862 0.801 0.817 0.513
-BB. Seg Acc 0.902 0.902 0.867 0.867 0.601
-EA. Angle Error 7.416 7.173 8.391 8.267 57.147
-Fit Ext. 0.0771 0.0527 0.0791 0.0741 0.1701
-Table 4: Comparison with Point2Cyl on DeepCAD dataset.
-Table 4 demonstrates the quantitative comparison.
+In Table 5(a), we present the quantitative comparison between our method and the most related approach, Point2Cyl.
 Compared with Point2Cyl, the metrics of the extrusion (Ext.) and barrel/based (BB.) segmentation are improved by $5.6\%$ and $3.5\%$, respectively, leading to $13\%$ and $32\%$ reduction on the extrusion axis (EA.) angle and fitting (Fit Ext.) error, respectively.
 This shows that by supervising the extrusion segmentation with curve loss (type and parameter loss), more geometrical information is introduced, leading to better geometry accuracy.
-We visualized the sketch predicted by Point2Cyl and Point2Primitive, as shown in Figure 6 (b).
+We visualized the sketch predicted by Point2Cyl and our method, as shown in Figure 5(b).
 It can be seen that because Point2Cyl represents the sketch by SDF, the edges of the sketch profiles are curved, while the counterpart of Point2Primitive  is sharp and accurate.
-Further, we present more visualizations of the predicted sketch by Point2Primitive in the Appendix 2.
-Figure 6 (c) demonstrates the segmentation results by the two methods.
-It can be seen that even though the multiple parts of one extrusion are labeled with different extrusion labels (rows 4 and 5), the segmentation network can still learn the geometry representation and cluster the points as the same extrusion.
-This implies that Point2Primitive can predict simplified primitives that are closer to human design.
+In addition, we present more visualizations of the predicted sketch by Point2Primitive  in the Appendix C.
+Figure 5(c) (b) demonstrates the segmentation results by the two methods.
+It can be seen that even though the multiple parts of one extrusion are labeled with different extrusion labels, the segmentation network can still learn the geometry representation and cluster the points as the same extrusion.
+This implies that our method can predict simplified primitives that are closer to human design.
 4.6 Ablation Study
-We perform ablation studies to carefully analyze the DeNoising (DN), the improved transformer decoder, the sketch primitive definition, and the balancing weight in the loss function.
+We perform ablation studies to carefully analyze the query denoising (DN), the improved transformer decoder, the sketch primitive definition, and the balancing weight in the loss function.
 All quantitative metrics are measured on the DeepCAD dataset.
-No. $\beta$ SkhRep ImpDec $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$
+(a) Accuracy curve.
+No. $\beta$ SkhRep ImpDec $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD
 0 2.0 - - 82.19% 72.49% 0.819
@@
 Ours 2.0 96.14% 86.81% 0.312
-Table 5: Ablation study. left panel is the accuracy curves during training; right panel is the quantitative results.
-DeNoising.
-It can be seen from the convergence curve shown in Figure 5 that by employing DeNoising in the training procedure, the ultimate sketch prediction accuracy is improved by $11.4\%$ and $3.1\%$ on $Acc^{SKH}_{t}$ and $Acc^{SKH}_{p}$, respectively.
-More validation loss curves during training are presented in Appendix 5.
-From the curves, we can see that by directly predicting the curve parameters and types, the accuracy is extremely high (type error below 0.5% and parameter error below 0.06)
+(b) Quantitative results of ablation study.
+Figure 6: Ablation study.
+Query denoising.
+It can be seen from the convergence curve shown in Figure 6(a) that by employing denoising in the training procedure, the ultimate sketch prediction accuracy is improved by $11.4\%$ and $3.1\%$ on $Acc^{SKH}_{t}$ and $Acc^{SKH}_{p}$, respectively.
+Please refer to Appendix.
+D for more visualizations.
+From the curves, one can see that by directly predicting the curve parameters and types, the accuracy is extremely high (type error below 0.5% and parameter error below 0.06).
 Decoder and Sketch Hierarchy.
 We replace the improved decoder layer (ImpDec) with the vanilla transformer.
-Also, we utilize the sketch representation (SkhRep) method in [37], which converts parameter prediction into a classification problem.
-As shown in Table 5, the $Acc^{SKH}_{t}$ improves by $12.2\%$ by utilizing the ImpDec.
+Also, we utilize the sketch representation (SkhRep) method in Wu_Xiao_Zheng_2021, which converts parameter prediction into a classification problem.
+As shown in Figure 6(b), the $Acc^{SKH}_{t}$ improves by $12.2\%$ by utilizing the ImpDec.
 Also, the center-prior sketch hierarchy contributes to an $8.4\%$ improvement in $Acc_{p}$.
 This shows the effectiveness of the proposed improved transformer decoder.
-Balancing Weight. Table 5 shows that the balance weight $\beta=2$ performs slightly better than the other conditions, and we set it to our default value.
+Balancing Weight. Figure 6(a) shows that the balance weight $\beta=2$ performs slightly better than the other conditions, and we set it to our default value.
 5 Conclusion
@@
 In further work, we plan to extend our approach to include more complex modeling commands.
-Supplementary Material
-1 Evaluation Metrics
-1.1 Primitive Type and Parameter Metrics
+Appendix A The Use of Large Language Models (LLMs)
+In this work, Large Language Models (LLMs) were used solely to polish the language for clarity and readability.
+No LLMs were employed for idea generation, experimental design, data analysis, or any other part of the research process.
+Appendix B Evaluation Metrics
+B.1 Primitive Type and Parameter Metrics
 The command type and command parameter accuracy evaluate how accurate the predicted command is. For each CAD command, we calculate the command type accuracy by:
-$ACC_{type}=\frac{1}{N_{c}}\sum\limits_{i=1}^{N_{c}}\Psi[t_{i}=\hat{t}_{i}]$ (7)
+$ACC_{\text{type}}=\frac{1}{N_{c}}\sum\limits_{i=1}^{N_{c}}\Psi[t_{i}=\hat{t}_{i}],$ (9)
 where $N_{c}$ denote the total number of the CAD commands, $C_{i}$ and $\hat{C}_{i}$ are the ground truth command type and predicted command type, respectively.
@@
 The command parameter accuracy is calculated by:
-$ACC_{param}=\frac{1}{K}\sum\limits_{i=1}^{N_{c}}\sum\limits_{j=1}^{\lvert\hat{p}_{i}\rvert}\Phi[\lvert p_{i,j}-\hat{p}_{i,j}\rvert<\epsilon]\Psi[C_{i}=\hat{C}_{i}]$ (8)
+$ACC_{\text{param}}=\frac{1}{K}\sum\limits_{i=1}^{N_{c}}\sum\limits_{j=1}^{\lvert\hat{p}_{i}\rvert}\Phi[\lvert p_{i,j}-\hat{p}_{i,j}\rvert<\epsilon]\Psi[C_{i}=\hat{C}_{i}],$ (10)
 where $K=\sum_{i=1}^{N_{c}}\Psi[C_{i}=\hat{C}_{i}]\lvert p_{i}\rvert$ is the total number of parameters in all correctly recovered commands.
@@
 In practice, we use $\epsilon=0.01$.
-1.2 Extrusion Segmentation IoU
+B.2 Extrusion Segmentation IoU
 To evaluate the extrusion segmentation, we use the Segmentation (Seg) IoU as the metric.
@@
 The Seg IoU can be formulated as follows:
-$\mathrm{Seg\ IoU}$ $=\frac{1}{K}\sum_{k=1}^{K}RIoU(\mathbbm{1}(\mathbf{\hat{W}}_{:,k}),\mathbf{W}_{:,k})$ (9)
-$RIoU(\mathbf{X},\mathbf{Y})$ $=\frac{\mathbf{X}^{\top}\cdot{\mathbf{Y}}}{{\left\|\mathbf{X}\right\|}_{1}+{\left\|\mathbf{Y}\right\|}_{1}-\mathbf{X}^{\top}\cdot{\mathbf{Y}}}$ (10)
+$\mathrm{Seg\ IoU}$ $=\frac{1}{K}\sum_{k=1}^{K}RIoU(\mathbbm{1}(\mathbf{\hat{W}}_{:,k}),\mathbf{W}_{:,k}),$ (11)
+$RIoU(\mathbf{X},\mathbf{Y})$ $=\frac{\mathbf{X}^{\top}\cdot{\mathbf{Y}}}{{\left\|\mathbf{X}\right\|}_{1}+{\left\|\mathbf{Y}\right\|}_{1}-\mathbf{X}^{\top}\cdot{\mathbf{Y}}},$ (12)
 where $\mathbf{\hat{W}}_{:,k}$ and $\mathbf{W}_{:,k}$ are the predicted and ground-truth labels of the $k$-th extrusion, respectively. $\mathbbm{1}(\cdot)$ indicates the one-hot conversion.
-1.3 Base/Barrel Point Segmentation
+B.3 Base/Barrel Point Segmentation
 The Base/Barrel (BB) point Segmentation accuracy is defined as follows:
-$\mathrm{BB\ Acc}=\frac{1}{N}\sum_{i=1}^{N}(\mathbbm{1}(\mathbf{\hat{B}}_{i,:})==\mathbf{B}_{i,:})$ (11)
+$\mathrm{BB\ Acc}=\frac{1}{N}\sum_{i=1}^{N}(\mathbbm{1}(\mathbf{\hat{B}}_{i,:})==\mathbf{B}_{i,:}),$ (13)
 where $\mathbf{\hat{B}}_{i,:}$ and $\mathbf{B}_{i,:}$ are the predicted and ground-truth barrel label, respectively. $N$ is the number of input point clouds.
-1.4 Extrusion Axis Angle Error
+B.4 Extrusion Axis Angle Error
 Extrusion axis (EA) angle error measures the angle error between the GT and prediction, which is defined as follows:
-$\mathrm{EA\ Angle\ Error}=\frac{1}{K}\sum_{k=1}^{K}\arccos(\mathbf{\hat{e}}^{\top}_{k,:}\mathbf{e}_{k,:})$ (12)
+$\mathrm{EA\ Angle\ Error}=\frac{1}{K}\sum_{k=1}^{K}\arccos(\mathbf{\hat{e}}^{\top}_{k,:}\mathbf{e}_{k,:}),$ (14)
 where $\mathbf{\hat{e}}^{\top}_{k,:}$ and $\mathbf{e}_{k,:}$ are the predicted and ground-truth extrusion axis, respectively.
-1.5 Extrusion Fitting Error
+B.5 Extrusion Fitting Error
 The extrusion fitting (Fit Ext) error measures the average fitting error of each extrusion, which can be defined as follows:
-$\mathrm{Fit\ Ext}$ $=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F}$ (13)
-$\mathcal{F}$ $\triangleq\frac{1}{N_{\mathbf{P}_{k}}}\sum\left|SDF(\hat{\mathbf{s}}_{k},\hat{\mathbf{S}}_{k})-SDF(\mathbf{s}_{k},\mathbf{S}_{k}))\right|$ (14)
-$\hat{\mathbf{S}}_{k}$ $=\prod\limits(\mathbf{P}^{barrel_{k}},\hat{e}_{k,:},\hat{c}_{k,:})$ (15)
+$\mathrm{Fit\ Ext}$ $=\frac{1}{K}\sum_{k=1}^{K}\mathcal{F},\quad\text{where}$ (15)
+$\mathcal{F}\triangleq\frac{1}{N_{\mathbf{P}_{k}}}\sum$ $\left|\texttt{SDF}(\hat{\mathbf{s}}_{k},\hat{\mathbf{S}}_{k})-\texttt{SDF}(\mathbf{s}_{k},\mathbf{S}_{k}))\right|,\quad\text{and}\quad\hat{\mathbf{S}}_{k}=\prod\limits(\mathbf{P}^{barrel_{k}},\hat{e}_{k,:},\hat{c}_{k,:}),$ (16)
 where $\hat{c}_{k,:}$ is the predicted extrusion center.
-($\prod$($\cdot$)∏ ( ⋅ ) projects per extrusion barrel points using the extrusion axis and center.
+$\prod(\cdot)$ projects per extrusion barrel points using the extrusion axis and center.
 The inner summation $\mathcal{F}(\mathbf{P},k)$ represents the goodness of the $k$-th extrusion fitting.
-1.6 Primitive Number
+B.6 Primitive Number
 The primitive number is a metric of reconstructed sequence fidelity.
 We calculate the $\Delta\#P$ to measure the length difference between the reconstructed and ground-truth sequence, which is defined as follows:
-$\Delta\#P=\left|\hat{N}_{c}-N_{c}\right|$ (16)
+$\Delta\#P=\left|\hat{N}_{c}-N_{c}\right|,$ (17)
 where $\hat{N}_{c}$ and $N_{c}$ are the predicted and ground-truth primitive numbers.
-2 Visualized Sketch Prediction
+Appendix C Visualized Sketch Prediction
 Figure 7: Visualization of the sketch primitive prediction.
@@
 However, some parameter errors can be seen in the Figure.
-Therefore, we develop a post-optimized method to eliminate the parameter errors, which is detailed in Section 5.2.
-3 Validation Loss w/o DeNoising
-Figure 8: Validation curves with and without DeNoising of the Point2Primitive during training: (a) the class error during training; (b) the L-1 loss of the primitive parameter during training; (c) the distance loss of the primitive parameter during training.
+Therefore, we develop a post-optimized method to eliminate the parameter errors, which is detailed in Section F.2.
+Appendix D Validation Loss w/o query denoising
+Figure 8: Validation curves with and without query denoising of the Point2Primitive during training: (a) the class error during training; (b) the L-1 loss of the primitive parameter during training; (c) the distance loss of the primitive parameter during training.
 Except for the command type and accuracy curve, Fig.
@@
 Also, the convergence time trained with DN is shorter than without DN.
-4 The Implementation of the Extrusion Segmentation Network
+Appendix E The Implementation of the Extrusion Segmentation Network
 Figure 9: The implementation of the extrusion segmentation network: (a) the network architecture of the Point2Primitive; (b) the Sketch Point Encoder based on the PointCNN
-4.1 Extrusion and Barrel Point Segmentaion
+E.1 Extrusion and Barrel Point Segmentaion
 Figure 9 demonstrates the extrusion segmentation network.
@@
 Thus the extrusion $\mathbf{\hat{W}}$ and barrel $\mathbf{\hat{B}}$ segmentation logits are formulated as follows:
-$\mathbf{\hat{W}}_{:,j}$ $=\mathbf{\hat{M}}_{:,2j}+\mathbf{\hat{M}}_{:,2j+1}$ (17)
-$\mathbf{\hat{B}}_{:,0}$ $=\sum_{i}\mathbf{\hat{M}}_{:,2j}$ (18)
-$\mathbf{\hat{B}}_{:,1}$ $=\sum_{i}\mathbf{\hat{M}}_{:,2j+1}$ (19)
-4.2 Sketch Point Encoder
-As shown in Figure 9, the sketch point encoder is developed based on the PointCNN[16].
+$\hat{\mathbf{W}}_{:,j}$ $=\hat{\mathbf{M}}_{:,2j}+\hat{\mathbf{M}}_{:,2j+1},$ (18)
+$\hat{\mathbf{B}}_{:,0}$ $=\sum_{j}\hat{\mathbf{M}}_{:,2j},\quad\text{and}\quad\hat{\mathbf{B}}_{:,1}=\sum_{j}\hat{\mathbf{M}}_{:,2j+1}.$ (19)
+E.2 Sketch Point Encoder
+As shown in Figure 9, the sketch point encoder is developed based on the PointCNN(li2018pointcnnconvolutionmathcalxtransformedpoints).
 Instead of using random or fps points, fixed points are set as the representation points.
 Also, to provide position encodings for the transformer, a learned position encoding is utilized.
-5 Post Optimization
-5.1 The Batch SDF Computing
+Appendix F Post Optimization
+F.1 The Batch SDF Computing
 We develop a batch SDF computing method to calculate the sketch SDF considering batch input.
@@
 We now provide an anonymous GitHub repository to present the SDF computing in https://github.com/AnonymousRepo1234/Point2Primitive/blob/main/SDF_batch_cal.py.
-5.2 Curve Fine-tuning
+F.2 Curve Fine-tuning
 Figure 10: Post optimization of curve prediction.
@@
 Algorithm 1 Curve Fine-tuning
-$Curves$
+CurvesCurves
 $T$ as modified Curves
 $Curves_{norm}\leftarrow normalize(Curves)$
-$Curves_{norm}\leftarrow sort(Curves_{norm})$triangleright sorted by starting point
+$Curves_{norm}\leftarrow sort(Curves_{norm})$ $\triangleright$ sorted by starting point
 $S\leftarrow Curves_{norm}$
-$T\leftarrow\emptyset$triangleright as modified Curves
+$T\leftarrow\emptyset$ $\triangleright$ as modified Curves
 while $S\neq\emptyset$ do
@@
      $T\leftarrow T\cup\{x\}$
-                    $S\leftarrow S\backslash\{x\}$
+     $S\{$$\leftarrow S\backslash\{x\}$
      break
@@
      $T\leftarrow T\cup\{x\}$
-                    $S\leftarrow S\backslash\{x\}$
+     $S\{$$\leftarrow S\backslash\{x\}$
      break
@@
 $T\leftarrow post\_process(T)$
-5.3 Loop Opmitzation
+F.3 Loop Opmitzation
 Algorithm 2 Loop Opmitzation
-$point\_cloud,Curves$
+point_cloud, Curves
 params as modified sketch curve params
-$eval\_points\leftarrow in\_pc(point\_cloud)$triangleright as pc to be formulated
-$cmd,params\leftarrow Initialize(sketch)$triangleright cmd as mask of curve type, params as shape of sketch
-for each epoch in $epoches$ do
-       $line\_loss\leftarrow SDF(eval\_points,params[\newline line\_mask(cmd)])$triangleright calculate line, arc and circle loss
-       $arc\_loss\leftarrow SDF(eval\_points,params[\newline arc\_mask(cmd)])$
-       $circle\_loss\leftarrow SDF(eval\_points,params[\newline circle\_mask(cmd)])$
-       $loss\leftarrow concat(line\_loss,arc\_loss,circle\_loss,-1)$
-       $loss.min()$triangleright minimize the last element of loss tuple
+eval_points $\leftarrow$ in_pc(point_cloud) $\triangleright$ as pc to be formulated
+cmd, params $\leftarrow$ Initialize(sketch) $\triangleright$ cmd as mask of curve type, params as shape of sketch
+for each epoch in epochesepoches do
+  line_loss $\leftarrow$ SDF(eval_points, params[line_mask(cmd)]) $\triangleright$ calculate line, arc and circle loss
+  arc_loss $\leftarrow$ SDF(eval_points, params[arc_mask(cmd)])
+  circle_loss $\leftarrow$ SDF(eval_points, params[circle_mask(cmd)])
+  loss $\leftarrow$ concat(line_loss, arc_loss, circle_loss, -1)
+  loss.min() $\triangleright$ minimize the last element of loss tuple
   loss.backward()
@@
 To connect the predicted curves into loops, we use an optimization method based on Greedy Algorithm as shown in Algorithm 2.
-6 Extrusion Parameters Calculation
-We calculate the parameters of the extrusion operation by the obtained barrel points.
-The parameters of the extrusion operation $E_{i}=(\hat{e_{i}},\hat{t_{i}})$ consists of the extrusion axis $\hat{e_{i}}$, extrusion position $\hat{o_{i}}$, and extrusion extent $\hat{t_{i}}$.
-Firstly, the optimal extrusion axis of an input extrusion cylinder PC is given by $\hat{e_{i}}=\mbox{argmin}_{e_{i},\|e_{i}\|=1}(e_{i}^{T}H_{\phi}e_{i})$, where:
-$H_{\phi}=\textbf{N}^{\top}\Phi_{barr}^{\top}\Phi_{barr}\textbf{N}-\textbf{N}^{\top}\Phi_{base}^{\top}\Phi_{base}\textbf{N}$ (20)
+Appendix G Extrusion Parameters Calculation
+[…]
+$H_{\phi}=\textbf{N}^{\top}\Phi_{barr}^{\top}\Phi_{barr}\textbf{N}-\textbf{N}^{\top}\Phi_{base}^{\top}\Phi_{base}\textbf{N},$ (20)
 where N denotes the normals of the input points.
@@
 The extrusion extent $\hat{t}$ is the point projection range on the extrusion axis, which can be calculated by Equation 21.
-$\begin{split}&\hat{t}_{min}=\min\limits_{P_{i}\in{P_{barr}}}(e_{i}\cdot(P_{i}-\hat{o}))\\ &\hat{t}_{max}=\max\limits_{P_{i}\in{P_{barr}}}(e_{i}\cdot(P_{i}-\hat{o}))\\ \end{split}$ (21)
-7 More Details of the Improved Transformer Decoder
-7.1 Command Parameter Noise
+[…]
+Appendix H More Details of the Improved Transformer Decoder
+H.1 Command Parameter Noise
 For each input training pair $(p_{i},S_{i})$, we add random noises to both their curve types and curve parameters.
@@
 For curve type noising, the GT command types are randomly flipped.
-7.2 Attention Mask
+H.2 Attention Mask
 An attention mask is used to prevent information leakage.
@@
 Given a sketch contains $N_{c}$ curves. The noised ground-truth curves of all the primitives are first divided into $K$ groups. The denoising part is then denoted as:
-$\begin{split}&q=\{g_{0},g_{1},...,g_{K-1}\}\\ &g_{j}=\{q_{0}^{k},q_{1}^{k},...,q_{M-1}^{k}\}\end{split}$ (22)
-where $g_{k}$ is the $-k\mbox{-}$th denosing group.
-$q_{m}^{k}$ is the $-m\mbox{-}$th query in the denoising group.
+$\begin{split}q=\{g_{0},g_{1},...,g_{K-1}\},\quad\text{and}\quad g_{j}=\{q_{0}^{k},q_{1}^{k},...,q_{M-1}^{k}\}\end{split}$ (22)
+where $g_{k}$ is the $k\mbox{-}$th denosing group.
+$q_{m}^{k}$ is the $m\mbox{-}$th query in the denoising group.
 Each denoising group contains $M$ queries where $M$ is the number of commands in one input PC.
 The attention mask $A=[a_{ij}]_{W\times{W}}$ can then be formulated as:
-$a_{ij}=\left\{\begin{aligned} 1&,\text{ if }j<K\times{M}\text{ and }\lfloor\frac{i}{M}\rfloor\neq\lfloor\frac{j}{M}\rfloor;\\ 1&,\text{ if }j<K\times{M}\text{ and }i\geq K\times{M};\\ 0&,\text{ else }\\ \end{aligned}\right.$ (23)
-where $a_{ij}=1$ means the $-i\mbox{-}$th query cannot see the $-j\mbox{-}$th query and $a_{ij}=0$ other wise.
+[…]
+where $a_{ij}=1$ means the $i\mbox{-}$th query cannot see the $j\mbox{-}$th query and $a_{ij}=0$ other wise.
 The decoder embeddings are taken as curve embeddings in our model.
-Therefore, an indicator is appended to the curve embeddings to distinguish the denoising part from the matching part, as shown in Figure 3, where $1$ means the query belongs to the denoising part and 00 means the query belongs to the matching part.
-8 Applications
+Therefore, an indicator is appended to the curve embeddings to distinguish the denoising part from the matching part, as shown in Figure 2, where $1$ means the query belongs to the denoising part and 0 means the query belongs to the matching part.
+Appendix I More visualizations
+We provide more visualization results as shown in Figure 11 and Figure 12.
+Figure 11: More visual comparison between reconstruction results on the DeepCAD and Fusion360 gallery dataset.
+Figure 12: More visual comparison between reconstruction results on the augmented DeepCAD dataset.
+Appendix J Applications
 Figure 13: Examples of some practical application: (a) reconstruction of the structural components; (b) reconstruction of the buildings; (c) reconstruction of the small urban agglomeration.